### v1

# Large Language Model powered Symbolic Execution

Abstract.
Large Language Models (LLMs) have emerged as a promising alternative to traditional static program analysis methods, such as symbolic execution, offering the ability to reason over code directly without relying on theorem provers or SMT solvers. However, LLMs are also inherently probabilistic by nature, and therefore face significant challenges in relation to the accuracy and scale of the analysis in real-world application. Such issues often necessitate the use of larger LLMs with higher token limits, but this requires enterprise-grade hardware (GPUs) and thus limits accessibility for many users.
In this paper, we propose LLM-based symbolic execution—a novel approach that enhances LLM inference via a path-based decomposition of the program analysis tasks into smaller (more tractable) sub-tasks. The core idea is to generalize path constraints using a generic code-based representation that the LLM can directly reason over, and without translation into another (less-expressive) formal language. We implement our approach in the form of AutoExe, an LLM-based symbolic execution engine that is lightweight and language-agnostic, making it a practical tool for analyzing code that is challenging for traditional approaches. We show that AutoExe can improve both the accuracy and scale of LLM-based program analysis, especially for smaller LLMs that can run on consumer grade hardware.
1. Introduction
Program analysis is a foundational discipline in computer science that aims to understand program behavior through various systematic techniques. Traditional forms of program analysis include static methods—such as symbolic execution (King, 1976), abstract interpretation (Cousot and Cousot, 1977, 1979), and model checking (Clarke and Emerson, 1982)—that analyze the program without executing it, as well as dynamic methods—such as fuzzing (Miller et al., 1990; Zeller et al., 2024), concolic execution (Godefroid et al., 2005; Sen et al., 2005), instrumentation (Serebryany et al., 2012; Nethercote and Seward, 2007), and profiling (Graham et al., 1982)—that analyze program behavior based on observations of actual executions. In general, program analysis has many applications, such as program testing (Hamlet, 1977; Clarke, 1976; Rapps and Weyuker, 1985; Zhu et al., 1997; Godefroid et al., 2005; Cadar et al., 2008; Sen et al., 2005; Pacheco and Ernst, 2007; Weyuker, 1983; Goodenough and Gerhart, 1975), debugging (Zeller, 2005; Agrawal et al., 1993; Engler et al., 2001), verification (Floyd, 1967; Hoare, 1969), repair (Le Goues et al., 2012; Long and Rinard, 2016), reverse engineering (Nelson, 2005), and vulnerability detection (Kaur and Nayyar, 2020). One recent alternative to traditional program analysis methods has emerged in the form of Large Language Models (LLMs) (Chen et al., 2021; Li et al., 2022; Touvron et al., 2023; Bubeck et al., 2023; OpenAI, 2024). Here, LLMs can make inferences over code directly with properties expressed in code or natural language and have become powerful enough to handle many traditional program analysis applications. As such, many LLM-based testing (Chen et al., 2024a; Meng et al., 2024c), debugging (Majdoub and Ben Charrada, 2024), and repair (Zhou et al., 2024a, b; Xia et al., 2023) tools have recently emerged.
We contrast traditional program analysis (symbolic execution) with LLM-based program analysis. Here, symbolic execution is a static program analysis method based on the idea of executing a program with symbolic inputs/values, typically represented as logical formulae over some underlying theories (e.g., linear arithmetic, bit vectors, and arrays). Symbolic execution systematically transforms these symbolic values (a.k.a. the symbolic state) based on the program’s statements, effectively executing multiple concrete paths simultaneously. These symbolic states can then be used for various analysis tasks, such as verifying the truth of an assertion or generating test cases, using deductive inference with the help of some underlying theorem prover or SMT solver (de Moura and Bjørner, 2008). In contrast, LLM-based program analysis involves engineering a prompt (or sequence of prompts) that queries an LLM as an oracle. A typical prompt consists of the relevant code (or code fragments), as well as instructions explaining the analysis task expressed in natural language. The LLM uses probabilistic inference (based on the training data) to solve the task, rather than the strict deductive inference used by a theorem prover. As such, LLM-based program analysis is typically ad hoc and tailored to specific tasks rather than general and principled.
Both traditional and LLM-based program analysis face significant challenges in practice. For example, traditional symbolic execution has several well-known limitations (Baldoni et al., 2018; Cadar and Sen, 2013; Meng et al., 2024a), such as: the handling of unbounded loops, the handling of external environment/libraries, and the handling of memory/heap-manipulating programs—all of which are common-place in real-world code. We consider how KLEE (Cadar et al., 2008), a prominent symbolic execution engine, treats each iteration of a loop as a separate path, leading to non-termination for unbounded loops (e.g., ). In contrast, LLMs have the ability to directly reason over loops, environment, and heaps, avoiding problems such as non-termination. That said, LLMs face other challenges, such as scalability (e.g., the 8192 token limit for GPT-4 (Bubeck et al., 2023)) and accuracy (Levy et al., 2024; Fang et al., 2024) due to probabilistic reasoning. Recent studies also correlate the accuracy of LLMs with prompt size (Levy et al., 2024), meaning that more concise and targeted prompts generally perform better.
In this paper, our aim is to improve the accuracy and scale of LLM-based program analysis. Our first main insight is that the strengths/weaknesses of traditional and LLM-based program analysis are complementary, meaning that a hybrid design can help address the limitations of either approach. To this end, we propose combining the path-based decomposition of symbolic execution with probabilistic inference via an LLM—a.k.a. LLM-based symbolic execution. The core idea is a principled decomposition of the original program analysis task into smaller sub-tasks based on paths in the original program—helping to mitigate some of the scalability and accuracy concerns with LLMs. Our second main insight is that since LLMs are primarily trained on code, the representation of symbolic states should also be in terms of code rather than logical formulae. To this end, we propose a generic path constraint representation of the form of a strongest post-condition () predicate transformer (Dijkstra, 1975) over sub-programs derived from the original program, where each sub-program represents a path. Since we represent path constraints as ordinary code, we can use an LLM prompting directly, disposing of any verification conditions (VCs) generated by the symbolic execution process. Our approach also avoids many of the limitations associated with the translation of paths into formulae for a theorem prover (e.g., environment and heaps.). Our final insight is that our path constraint representation can be generalized into sets of paths (e.g., all iterations of an unbounded loop), ensuring that LLM-based symbolic execution will always terminate.
We study LLM-based symbolic execution both in theory and practice. We show that LLM-based symbolic execution mitigates many of the limitations of both traditional and LLM-based program analysis. We have implemented our approach in the form of AutoExe—an automated LLM-based symbolic execution engine. AutoExe uses a path-based decomposition of a program analysis task into smaller (more-tractable) sub-tasks that are suitable for LLM inference. Based on the observation that LLMs are inherently approximate oracles, we propose a lightweight design of AutoExe that is language agnostic and does not rely on any heavyweight compiler infrastructure. AutoExe is designed to be a practical program analysis tool that can be applied to code that is difficult-to-analyze with traditional methods. In summary, the main contributions of this paper are:
We introduce the concept of LLM-based symbolic execution—a symbolic execution methodology using LLMs for direct reasoning over the original programming language (e.g.,
C
/Java/Python/etc.). Our approach uses a path-based decomposition of program analysis tasks into more concise LLM prompts, improving the accuracy and scalability of LLM inference. -
We implement our approach in the form of AutoExe—a lightweight LLM-based symbolic execution engine that supports multiple programming languages without relying on heavyweight compiler infrastructure.
We evaluate AutoExe against various program analysis tasks for
C
/Python/Java code. We show that AutoExe improves both accuracy and scale, especially for smaller models that can run on consumer grade hardware.
2. Motivation
2.1. Background
2.1.1. Symbolic Execution
Symbolic execution (King, 1976; Baldoni et al., 2018) is an established program analysis methodology based on running the program with symbolic values representing sets of concrete inputs (rather than a single concrete input for normal execution). For each symbolic input, symbolic execution will systematically explore multiple execution paths of the program, allowing for applications such as error detection, verifying correctness, or finding vulnerabilities.
Symbolic execution works with symbolic states, a.k.a. path constraints, that are traditionally represented as a set of variables (e.g., x
, y
, etc.) subject to constraints (e.g., , etc.). The symbolic state represents the set of concrete states (e.g., ) that is reachable from the symbolic input. Symbolic execution can be defined in terms of operations over symbolic states, e.g., symbolic execution over C
-style increment statement (x++
) can be represented as the Hoare triple (Hoare, 1969):
Given a symbolic state (), the triple describes the resulting symbolic state () after the execution of the increment operator. For example, given , then , or equivalently, . Conditional statements (e.g., ) are typically handled by forking the symbolic state () into separate and , and then continuing the execution along the individual then- and else-branches. As similar strategy is used for loops (e.g., ). Finally, for each path () through the program, a final path constraint will be generated. This can be used to prove properties over the path, such as whether a final post-condition () holds, i.e., whether the verification condition (VC) of the form () holds or not. The VCs can be (dis)proved (i.e., disposed of) with the help of a suitable theorem prover, such as an SMT solver (de Moura and Bjørner, 2008). By executing sets of concrete inputs at once, symbolic execution will exhaustively explore the (infinite) space of program behaviors—something cannot be explored by concrete execution alone.
Example 2.0.
Consider the simple program () and the post-condition . Then the following VC will be generated for the -path:
(VC-Else) |
A theorem prover disproves (VC-Else), meaning () does not hold for the whole program. ∎
2.1.2. Large Language Models
The emergence of Large Language Models (LLMs) (Chen et al., 2021; Li et al., 2022; Touvron et al., 2023; Bubeck et al., 2023; OpenAI, 2024) presents a new alternative for reasoning over program code. Trained on extensive datasets, including millions of lines of real-world code written in multiple languages, LLMs can reason over C
/Java/Python/etc. code directly, including many traditional program analysis tasks, such as testing (Chen et al., 2024a; Meng et al., 2024c), debugging (Majdoub and Ben Charrada, 2024), repair (Zhou et al., 2024a, b; Xia et al., 2023), etc. For example, instead of relying on symbolic execution, LLMs can reason over code using a suitable prompt expressed in natural language, such as “there is a bug present in the following Python code segment, please suggest the possible root causes of the bug and corresponding fixes”. Most modern LLMs can analyze the code, and generate possible suggestions and patches automatically, based on the understanding of code and defects present in the training set.
Unlike traditional program analysis methods, LLMs do not aim to be perfectly precise. Rather, LLMs can be thought of as approximate oracles that are sometimes incomplete or give the wrong answer. This is because LLMs fundamentally rely on learned patterns and probabilistic reasoning, rather than classical deductive reasoning used by traditional theorem provers. Despite the difference, LLMs are clearly useful, with an explosion of applications for real-world analysis problems.
(a) original program
(b) slice for “mutx is unlocked”
(c) slice for “”
2.2. Limitations of Symbolic Execution
While symbolic execution has many applications (e.g., bug detection, security analysis, debugging, and program repair, etc.), it also has well-known limitations regarding loops, environment, and heap manipulating programs. We elaborate on the ideas from (Meng et al., 2024a), as summarized below:
2.2.1. Limitation: Handling Loops (and Recursion)
Unbounded loops (and recursion) are a known problem for symbolic execution. Here, symbolic state forking (Section 2.1.1) can lead to an infinite unrolling of the loop, where each possible loop iteration is treated as a separate path. Popular symbolic execution tools, such as KLEE (Cadar et al., 2008), handle this problem using a concrete iteration bound—exchanging potential infinite exploration with an incomplete exploration, as necessary.
An alternative is to use loop invariants. If known, the invariant allows symbolic execution to pass over a loop without explicit unrolling. Loop invariants can be manually provided, or discovered automatically, such as using abstract interpretation (Cousot and Cousot, 1977) over some known domain, constraint solving (CBMC (Kroening and Tautschnig, 2014)), or machine learning (Code2Inv (Si et al., 2018)). However, loop invariant discovery may only work for simple loops, and the general case is either computationally hard or undecidable.
2.2.2. Limitation: Handling External Environment
Another known problem is the handling of external functions (e.g., calls to third-party libraries without source code) and/or external inputs (e.g., recv
from a socket), collectively the external environment. Since the underlying symbolic execution theorem prover uses deductive reasoning, a precise specification of all external operations and/or inputs is usually required. As such, the environment is usually handled through a combination of stubs, modeling, or concretization. For example, the user can manually model an external function call by implementing a replacement stub function that specifies necessary specification using klee_assume()
. However, this approach is manual, and modeling arbitrary code or inputs can require significant effort, meaning that the approach tends to rarely scale.
Another approach is concretization, where the symbolic execution algorithm assigns concrete values to some symbolic variables, allowing external functions to be executed with these values. However, concretization can also lead to an incomplete exploration of program behavior.
2.2.3. Limitation: Heap Manipulating Programs
Traditional theorem provers and SMT solvers tend to have limited support for reasoning over (mutable) data-structures with complex structural invariants, such as singly- and doubly-linked lists, binary trees, red-black trees, DAGs, etc. By extension, traditional symbolic execution tools inherit these limitations. Some tools are based on Separation Logic (Reynolds, 2002), which does support reasoning over heap manipulating programs, as used by VeriFast (Jacobs et al., 2011) and Infer (Distefano et al., 2019). However, VeriFast is manual and annotation-heavy, and Infer is limited to heuristic-based inference based on common structural invariants, such as linked-lists.
Discussion
Figure 1 (a) is an example of a program that exhibits all three limitations, including unbounded loops (while (true) ...
), interaction with the external environment (recv
/send
), and is a heap manipulating program (NODE
). Although this program is relatively simple, it still presents a significant challenge for traditional symbolic execution tools such as KLEE.
2.3. Limitations of LLMs
LLMs are trained on a huge corpus of data, and do not necessarily have the same limitations of symbolic execution. For example, an LLM can reason over Figure 1 (a), and can answer queries about the program. That said, LLMs also have known limitations, as summarized below:
2.3.1. Limitation: Scale
LLMs typically have a limited ability to reason over large and complex code bases. For example, the token limit imposes a maximum number of tokens for the given input (and output), and a medium-to-large code base can easily exceed this limit.
2.3.2. Limitation: Approximate Oracles
LLMs are approximate oracles in that they use probabilistic reasoning, and may not necessarily generate accurate answers. Even if the token limit (2.3.1) is not exceeded, studies have shown that LLMs generally perform worse with overly verbose prompts that include irrelevant information (Levy et al., 2024). This concept could be described as a soft token limit, i.e., that prompts should be concise and targeted in order to maximize LLM accuracy.
Discussion
Although scale and accuracy are concerns, LLMs can typically handle more classes of programs than traditional analysis methods such as symbolic execution. Furthermore, the relative completeness of LLMs, in the absence of precise specifications, means that LLMs can be easily applied to a wide range of applications. The completeness/applicability can often be of greater pragmatic interest than perfect accuracy, and this is one reason behind the explosion of real-world applications.
2.4. Our Approach
Many of the limitations of traditional symbolic execution are limitations of the underlying theorem prover. Specifically, existing theorem provers and SMT solvers only accept queries in some formal input language, with limited expressiveness compared to the original source language (e.g., C
/Java/Python). For example, an SMT solver will only accept queries in the form of a Boolean formula over a given set of theories (T), such as linear inequalities, bit vectors, or arrays. In contrast, the C
programming language is significantly more expressive, with complex control-flow (loops), memory, pointers, library calls, data-structures, environmental interactions, etc. The discrepancy in the expressiveness complicates the translation from the high-level programming language into the solver input language—such as “unrolling” loops into flat quantifier-free formula—and such translation may also be incomplete (e.g., unbounded loops). Furthermore, the SMT solver may not support the necessary theories for reasoning over complex programs, such as heap manipulating data-structures, external environment interactions, etc.
Our underlying approach is to use the path-based decomposition of symbolic execution, but to replace the traditional theorem prover with an LLM. The key advantage of LLMs is that they can reason over the source code directly—eliminating the need for translation into a less expressive solver input language and the associated limitations. Instead, our approach represents the path constraint as a generic strongest post-condition predicate transformer (Dijkstra, 1975) over a pre-condition , and a derived sub-program which represents a path or set of paths. Our key insight is that LLMs can reason over -constraints directly, since is just ordinary source code, without the need for “translation”. Essentially, we can view the LLM as an effective solver for untranslated -constraints “as-is”. We demonstrate this concept with a simple example.
Example 2.0 (Simple).
Consider the simple program () and post-condition once more (see Example 2.1). Then the path constraint for the -branch can be represented as a formula () or a sub-program (), as follows:
Formula () | Strongest Post-Condition over a Sub-program () |
The key idea is that both () and are equivalent representations of the same path constraint—i.e., one can be derived from the other under the definition of the language semantics (). While the formula is suitable for a theorem prover, the sub-program is suitable for an LLM:
(VC-Else-2) | “Given , does the post-condition always hold?” |
The LLM determines the VC does not hold, e.g., the following GPT4 (Bubeck et al., 2023) output (emphasis original):
“The post-condition does not always hold. A simple counterexample is when , where instead of . Hence, the claim is false.”
Our approach is to enumerate all sub-programs based on a partitioning of paths through the original program. Here, each sub-program is algorithmically derived from the original code, and contains all statements that (1) are visited by any path from the partition, and (2) of which post-condition is data- or control-flow dependent. Like traditional symbolic execution, our approach is a path-based decomposition of the original program analysis problem into smaller (more tractable) sub-problems. This decomposition helps to address some of the limitations of direct LLM-based reasoning (Section 2.3), such as scale and accuracy. Furthermore, since our approach uses an LLM, it avoids many of the limitations of traditional symbolic execution (Section 2.2) caused by translation. We summarize the benefits as follows:
Handling Loops (and Recursion). Our approach avoids translation of (unbounded) into a less expressive language, and loops/recursion can be represented “as-is” in the derived sub-program(s). Similarly, our approach does not need explicit loop invariant recovery or annotation, as LLMs are capable of reasoning over loops without any special intervention.
Handling External Environment. Rather than manual modeling or concretization, our approach is to use the LLM to infer the likely behavior of the environment or external function call. Since LLMs are trained on a huge corpus of real-world code, they have significant exposure to common libraries, file formats, protocols, etc. Furthermore, even if the external environment is novel, LLMs can still infer the most likely behavior based on clues from the context (function names, variable names, code comments, placement within the algorithm, etc.), as a form of abductive reasoning, or inference to the best explanation, without the need for explicit modeling.
Heap Manipulating Programs. LLMs can directly interpret heap-manipulating programs without the need for any special logical framework. LLMs can also (abductively) infer the (likely) structural invariants based on direct interpretation of the code, without explicit annotation.
Scale. Like traditional symbolic execution, our approach decomposes program analysis problems into smaller (tractable) sub-problems, helping to avoid any hard or soft limit of the LLM. This allows our approach to scale to large/complex programs and analysis problems.
Approximate Oracles. Studies (Levy et al., 2024) show that LLMs perform better with more targeted and concise prompts. By decomposing program analysis problems into sub-problems that capture only the relevant parts of the original (possibly large) code base, we help to focus the LLM and improve the overall accuracy of the analysis.
The decomposition and lack of translation mean that our approach can handle programs that are difficult for traditional program analysis. We illustrate with an example.
Example 2.0 (LLM-based Symbolic Execution).
Consider the Figure 1 (a) program that cannot easily be handled by traditional symbolic execution methods (Section 2.2), and a simple program analysis problem that verifies each lock(&mutx)
operation is paired with an unlock(&mutx)
operation. Furthermore, assume that (for the sake of example) the (a) program is too complex for an LLM to handle directly (Section 2.3).111This is not necessarily true, but is an assumption for the sake of an example that can fit within the page limit. We can express this as a natural language pre- and post-condition ( and respectively) that “mutx
is unlocked”. Then Figure 1 (b) is an example of a derived sub-program that acts as a substitute for the original program with respect to and . We have that:
Thus, to prove for holds (a), it is sufficient to prove that holds for (b). Furthermore, program (b) is targeted to the specific program analysis task (that mutx
is unlocked), and is smaller than in terms of token count (419 vs 81).222As counted by https://platform.openai.com/tokenizer at the time of writing The LLM can determine that VeriCond holds using the program (b).
Another example is Figure 1 (c) and the post-condition (). Assuming the same condition initially holds, then (c) is smaller (419 vs 149 tokens). In this case, the LLM determines the post-condition does not hold using (c), since strdup()
may return NULL
Input: A Hoare triple Output: HOLDS or a counter-example 1 2 3Function LLMSymExe(): 4 5 6 7 for do 8 9 10 11 "does hold?" 12 13 if then return 14 15 return HOLDS 16 Algorithm 1 LLM-based Symbolic Execution
Algorithm
An overview of our LLM-based symbolic execution algorithm is summarized in Algorithm 1. Here, the algorithm’s frontend is similar to that of a standard compiler, and parses the program into an Abstract Syntax Tree (AST, line 2), and then generates a Control Flow Graph (CFG, line 3). Next, the algorithm generates a representation of the set of all paths through the CFG (line 4). Here, the set of all paths is represented as a set of path such that (), where each partition represents some (possibly infinite) subset of paths. One challenge is how to generate a good set of partitioning (to be discussed in Section 3). Next, for each partition , the algorithm generates a derived sub-program that generalizes the partition (line 6). For example, Figure 1 (b) and (c) are possible derived sub-programs of Figure 1 (a). Each sub-program is used to construct a corresponding prompt (line 7), including rendering the back into a text-based source-code representation (line 8). The prompt (line 7) is a natural language representation of the Hoare triple , which holds iff . Finally, the prompt is sent to the LLM for inference (line 10). Algorithm 1 can be used to prove or disprove the post-condition modulo the LLM’s reasoning capabilities. Assuming that the partitions are ordered based on slice, Algorithm 1 will return the least sub-program that is deemed to refute the post-condition. Otherwise, if no such refutation is found, Algorithm 1 deems that the triple holds (HOLDS
Property | Tradition Symbolic Execution | LLM-based Symbolic Execution | |
Design | Overall method | Decomposition & solving path constraints | Decomposition & solving path constraints |
Decomposition method | Decomposition into a formal language | Decomposition into sub-programs | |
Reasoning engine | Theorem prover or SMT solver | Large Language Model (LLM) | |
Reasoning method | Deductive | Probabilistic, deductive, abductive | |
Path representation | Formal language (unfolded -path constraints) | Untranslated -constraints over truncated slices | |
Specification language | Formal language | Either formal, code, or natural language | |
Cap. | Weaknesses | Loops, environment, heaps | Complex integer, linear, Boolean reasoning |
Unbounded loops | Infinite unfolding or loop invariants | LLM reasons over loops “as-is” | |
Environment | Manual modeling/specifications required | LLM abductive reasoning or training data | |
Heap manipulation | Manual annotation + Separation Logic | LLM reasons over heaps “as-is” | |
Imp. | Programming languages | Language specific (C+KLEE (Cadar et al., 2008), Java+SPF (Păsăreanu and Rungta, 2010)) | Programming language agnostic |
Compiler infrastructure | Close integration (LLVM (Lattner and Adve, 2004)+KLEE, etc.) | Lightweight (AST-level) implementation |
Summary. Like traditional symbolic execution, LLM-based symbolic execution (Algorithm 1) represents a path-based decomposition the original program analysis task into smaller sub-tasks. A summary of the main similarities and differences are shown in Figure 2. The substitution of a theorem prover with an LLM changes to various aspects of the design, capabilities, and implementation of the symbolic execution engine. For example, LLMs use probabilistic and adbuctive reasoning, or rely on information learned during the training process, meaning that LLMs do not need precise specifications or environment modeling. Likewise, the lack of translation into some (less expressive) formal language allows the LLM to reason over loops or heap manipulation “as-is”, without relying on loop/data-structure invariant discovery.
In this paper, we study the concept of program analysis via LLM-based symbolic execution. First, we study the principles of LLM-based symbolic execution in terms of the idealized procedural programming language used by Hoare logic (Hoare, 1969). We show that program analysis tasks can be decomposed into tasks over derived sub-programs representing paths, or sets of paths (truncated slices), through the original program. Furthermore, we also show that only a finite number of (Algorithm 1, line 4) needs to be considered, ensuring that LLM-based symbolic execution will always terminate—even for unbounded loops.
In addition, we study the application of LLM-based symbolic execution in practice. For this, we design AutoExe—an LLM-based symbolic execution engine for real-world programming languages such as C
/Java/Python. Our approach is based on the observation that Algorithm 1 is mostly language agnostic except for specific aspects, such as the parser. This means our approach can be readily ported to other programming languages. Furthermore, we also observe that, since LLMs are fundamentally approximate, we can build a lightweight implementation that uses approximate parsing and dependency analysis—without relying on any heavyweight and/or language-specific compiler framework.
3. Principles of LLM-based Symbolic Execution
Our goal is to adapt traditional symbolic execution methods to LLMs that reason over code directly, rather than translation into a (less expressive) theorem prover input language.
3.1. Symbolic Execution Foundations
We use the minimal imperative language defined by Hoare logic (Hoare, 1969) augmented with an explicit -statement.333Note can be emulated as () under partial correctness, but is treated as a special case. We define the language syntax as follows:
Where represents some base language (e.g., arithmetic expressions) and represents Boolean expressions over . We also use to sometimes represent an empty program (equivalent to ). The language semantics are defined inductively (i.e., least relation) in terms of the strongest post-condition () relation defined as follows:
We define a linear program to be any program comprising only -, -, and assignment-statements (without conditionals or loops). We define a to be a linear program that is derived by unfolding the original program using the following rules:
We abstractly define symbolic execution to be any algorithm that combines path unfolding with Verification Condition (VC) solving. Given a program analysis task represented as a Hoare (Hoare, 1969) triple , then a symbolic execution algorithm:
exhaustively generates the set of all through the program; and
solves each corresponding VC () for each .
That is, a symbolic execution algorithm computes the following using explicit enumeration:
The triple is deemed to hold if each individual VC holds for the corresponding path (), and to not hold otherwise. Here, each is defined to be the path condition for the corresponding path . Essentially, a symbolic execution algorithm decomposes the original program analysis task into simpler sub-tasks that can be solved separately.
3.1.1. Traditional Symbolic Execution
“Traditional” symbolic execution algorithms solve each VC using a suitable theorem prover, such as an SMT solver (de Moura and Bjørner, 2008). To do so, the -constraints for each path are translated into a logical formula over the domain of , by applying the (linear program subset of the) -rules defined above. The translated VC () is then solved using a theorem prover. Note that the translation is necessary since traditional theorem provers are limited to a specific input language (e.g., SMT-LIB), and cannot make inferences over an abstract -constraint directly. In addition, most practical symbolic execution tools implement several optimizations, including incremental path unfolding, incremental -translation, pruning infeasible paths, and the merging of similar paths. For example, rather than pre-computing the set of paths upfront, most practical implementations maintain a symbolic state comprising a current location, a (partially constructed) path constraint/condition, and a set of current variables-of-interest. Such optimizations are consistent with abstract symbolic execution (SymExe) defined above.
We can also understand the limitations of Section 2.2. Firstly, the number of paths in the set may be very large (exponential) or even infinite (unbounded loops, Section 2.2.1). This is known as the path explosion problem, and is a well-known limitation of symbolic execution methods. Secondly, the -translation will be limited for programs that contain operations that are not supported—such as common interactions with the external environment (Section 2.2.2). Finally, the underlying theorem prover itself may be limited. For example, KLEE (Cadar et al., 2008) uses the Z3 (Baldoni et al., 2018) SMT solver over the domain of linear arithmetic, arrays, and bit-vectors by default. However, this configuration does not support reasoning over heap manipulating programs (Section 2.2.3).
(a) simple program () | (b) path () | (c) truncation () |
1 i 2 while i n do 3 4 if x then 5 xs. 6 7 else 8 xs. 9 10 z 11 12 i 13 | 1 i 2 3 4 5 xs. 6 z 7 8 i 9 10 11 12 xs. 13 z 14 15 i 16 | 1 i 2 while i n do 3 4 5 xs. 6 z 7 8 i 9 (d) slice ( 1 i 2 while i n do 3 4 xs. 5 i 6 |
3.2. LLM-based Symbolic Execution (Path-based)
The basic idea behind “LLM-based” symbolic execution is to use a Large Language Model (LLM) as the underlying reasoning engine instead of a theorem prover. Thus, given a VC of the form (), we can use the LLM to reason directly over the path constraint and post-condition . This is possible since, through its training process, the LLM can interpret both the syntax of the path (represented as ordinary code), as well as the language semantics represented by the -rules. We illustrate with the following example.
Example 3.0 (Path-based LLM-based Symbolic Execution).
We consider the simple example shown in Figure 3 (a) which is based on (Korel and Laski, 1988) Figure 7 (a.k.a. Example Program 3). We assume that xs is a data-structure implementing a multi-set, and // can be expanded to sub-programs with a suitable multi-set implementation (e.g., a singly linked list), and is initially empty. Furthermore, we assume that the () operation always returns positive number, which can be expressed in natural language or as a formal rule (). Under these assumptions, the final size of xs should equal n. We can express this as the following triple:
Although conceptually simple, the Figure 3 (a) program is challenging for several reasons, including an unbounded loop, environmental input (), and data-structure reasoning (xs). We may prove (Triple) by enumerating paths, such as () from Figure 3 (b) representing two iterations of the loop. We can encode the VC () as a prompt, which is confirmed using an LLM:
“Given the pre-condition and the code , does the post-condition hold?” ∎
This example shows how LLMs can solve tasks that are challenging for traditional program analysis methods. That said, a simple path-based decomposition still inherits some limitations. Firstly, there can still be an infinite number of paths (i.e., path explosion), leading to non-termination. Secondly, each individual path could still be too long for the LLM to effectively reason over (e.g., path from Figure 3 (b) can be generalized to any length). Finally, paths can accumulate irrelevant statements, such as variable z, that can also exacerbate the path length problem.
3.3. LLM-based Symbolic Execution (Slice-based)
To address the path-explosion problem, our core idea is to merge individual -constraints (representing individual paths) into generalized -constraints representing (possibly infinite) sets of paths. Given a partition , our approach constructs a truncated sub-program such that:
Thus, instead of disposing of a (possibly infinite) number of verification conditions (VCs) of the form () for each , our approach disposes of a single VC of the form () for the entire set (). First, we shall present a method for constructing a truncated sub-program for a given partition. Next, we shall present a partitioning algorithm that need only consider a finite number of subsets, even for programs with infinite paths (unbounded loops), ensuring the symbolic execution algorithm always terminates.
3.3.1. Construction
We consider the construction of truncated sub-programs.
Truncation. Given a program and a (possibly infinite) subset of , we derive a truncated sub-program equivalent to for all , and is unreachable otherwise:
Here, () excludes all paths that terminate abnormally via a special (i.e., assume false) statement. Thus, represents a statement that is assumed to be unreachable.444Similar to __builtin_unreachable() from gcc. Consider all statements () in that are not covered by any path , then we can construct by replacing all such () with . We can represent this idea using the following rewrite rule:555We use the standard notation () to mean that any term matching the is rewritten to the term matching .
We may also apply the following rules to further simplify the resulting sub-program:
These rules preserve the (Truncation) property while also reducing the size (token count) of the resulting , which is beneficial for LLM prompting.
Slicing. We can further reduce the size of the truncated sub-program using program slicing (Weiser, 1981). We define a slice () to be a sub-program derived from by deleting (i.e., replacing by ) any statement () in that does not violate the condition:
Our main result is as follows. If () holds, then:
() holds by (Slice); and
() for all holds by (1) and (Generalization).
In other words, the single VC is sufficient to dispose of the entire partition (). We can apply standard slicing methods, such as Weiser’s back slicing algorithm (Weiser, 1981) (with as the slice criterion), to construct from . The back slicing algorithm traverses the Control Flow Graph (CFG), and deletes any statement that is not data- or control-flow-dependent on . The slicing algorithm is illustrated in Algorithm 2.
Example 3.0 (Slice-based Symbolic Execution).
We consider Example 3.1 once more. Here, the truncated slice () in Figure 3 (d) is a generalization of the infinite partition () representing all feasible paths through the -loop, including the Figure 3 (b) path (). An LLM can be used to verify the generalized verification condition (VC) encoded in natural language.
“Given the pre-condition and the code , does the post-condition hold?”
Path-based symbolic execution will infinitely unroll the -loop, generating a new VC for each loop iteration. In contrast, slice-based symbolic execution requires only a single VC to be checked. Furthermore, (, 5 lines, 28 tokens) is simpler and more compact than all of (, 10 lines, 56 tokens), (, 8 lines, 48 tokens), and (, 16 lines, 85 tokens). Truncation and slicing are useful for removing irrelevant statements while preserving paths, meaning that the corresponding prompt is simpler and more concise, thereby improving LLM accuracy. ∎
Discussion. Our approach is related to path merging and loop invariants in traditional symbolic execution. Here, given a set of path constraints, represented as translated logical formulae , the idea is to find a formula that is generalization (). Thus, the verification conditions () can be combined into a single verification condition (), helping to mitigate the path explosion problem. Similarly, the (possibly infinite) set of path constraints through a loop can be generalized into a loop invariant , such that (), allowing the loop to be handled without infinite unrolling. However, loop invariants discovery over formulae is difficult and undecidable in the general case. In contrast, our approach avoids the problem, since the -constraints are never translated into a different representation.
Input: A CFG sub-graph ; reverse topological order Output: A 1 2 3Function GenSlice(, ): 4 ; getVars() 5 for do 6 if modifies any , or 7 conditional reads any then 8 9 10 11 12 end for 13 return 14 15 16 Algorithm 2 Basic back-slicing algorithm.
3.3.2. Partitioning
Section 3.3.1 describes the construction of a truncated slice given a partition () that subset of paths (). In principle, any partitioning () can be used. However, a good partitioning aims to minimize the slice () size for each , , in order to simplify the prompts ultimately sent to the LLM. Our approach is to construct a partitioning based on path coverage.
Path Coverage. Under Section 3.1, we define a path () to be a linear program over statements or branches (represented as assertions) from . Here, we define a Control Flow Graph (CFG) path to be a sequence nodes (a.k.a. locations) through the CFG representation of . Given a CFG path (), we define path coverage to be the set (), i.e., the reinterpretation of () from a sequence to a set. It is common for distinct paths to have the same coverage, e.g., different iterations of the same loop with be distinct sequences, but will be equivalent sets (i.e., cover the same nodes). We make two key insights. First, paths with distinct coverage will have distinct truncations, since each path must differ by at least one location. This will be used as the basis for partitioning. Secondly, the number of distinct paths w.r.t. coverage is finite, meaning that LLM-based symbolic execution (slice-based) necessarily terminates—even for unbounded loops.
Input: The Control-Flow Graph (CFG) Output: A coverage-based partitioning Globals : Coverage map (), end node () 1 2 3Function GenPartitions(): 4 if then return 5 6 7 8 if then return 9 10 for do 11 12 GenPartitions() 13 return 14 Algorithm 3 Path partitioning algorithm
Partitioning Algorithm. The core idea is to enumerate CFG paths () with distinct coverage (). Each partition () is implicitly defined as all paths with the same path coverage as (). The algorithm uses a Depth First Search (DFS) exploration over the CFG, and is illustrated in Algorithm 3. Here, the algorithm takes a CFG representation fo , and generates a partitioning as a set of representative paths () for each distinct coverage class (). The algorithm works are as a standard DFS path construct algorithm, but also maintains a coverage map () that tracks previously-seen coverage prefixes. When a () is visited, the coverage map is consulted (line 2), and the path is pruned if the prefix has been observed before. The coverage map ensures that Algorithm 3 both (1) terminates, and (2) only returns paths (partitions) that differ by at least one location/node. Each output path () corresponds to a partition (), which is then used to construct a corresponding truncated sub-programs () and slice () which is used for LLM prompting.
Given two distinct coverages (, ), the resulting slices (i.e., , ) can sometimes be equivalent () if all distinguishing nodes are removed by the slicer. Furthermore, some truncated slices may generalize others, i.e., (). For example, under Algorithm 3, then is also a valid slice of (, Figure 3) representing zero iterations of the -loop. Nevertheless, our LLM-based symbolic execution algorithm orders VCs by size, in order to find a least refutation of the post-condition ().
Discussion. Algorithm 3 is guaranteed to terminate, since the number of possible distinct coverage sets is finite. Thus, unlike path-based symbolic execution, slice-based symbolic execution will always terminate, even for programs with unbounded loops. For example, the truncated slice () from Figure 3 (b) generalizes infinitely many paths through the loop (for any number of iterations). Furthermore, the size of each slice is always bounded by the size of the original program , whereas unrolled paths can exceed any length.
4. LLM-based Symbolic Execution in Practice
The truncated slice-based symbolic execution algorithm of Section 3 is defined for an idealized programming language (the language of Hoare triples). In this section, our aim is to port our abstract approach to real-world programming languages (e.g., C
/Java/Python). Our first main insight is that the main symbolic execution and slicing algorithms are agnostic to the target programming language—provided there is a method for parsing the source code into some suitable representation for path generation and slicing. Our second main insight is that, since LLMs are approximate oracles, the greater workflow can also be approximate—meaning that each individual step need not be perfectly precise, provided the overall accuracy is not significantly degraded. These insights significantly simplify our overall design.
An example of the workflow is illustrated in Figure 4. First, our workflow uses a lightweight parsing framework (specifically, tree-sitter
(Brunsfeld et al., 2025)) to parse the source code (a) into an Abstract Syntax Tree (AST) representation (b). From the AST, a Control-Flow Graph (CFG) is constructed, allowing for the finite enumeration of all path partitions using Algorithm 3. For each enumerated partition, a truncation and slice is constructed (c), by removing statements that are not control- or data-flow dependent on the post-condition , using Algorithm 2. Next, the slice is rendered back into the original source language (d). Essentially, this means emitting sliced path elements, deleting (i.e., not emitting) non-sliced path elements, and replacing (i.e., replace by assume(0)
) all other non-path elements. The resulting slice generalizes a (possibly infinite) set of paths through the program. Finally, the slice is used to construct a text-based prompt (e) used to query an LLM (f). steps (c)-(f) are repeated for each partition using Algorithm 1, until exhaustion or the discovery of a counter-example.
4.1. Partitioning and Truncated Slice Generation
Parsing.
We use the tree-sitter
framework (Brunsfeld et al., 2025) to parse the source code into an AST. Each AST node comprises a node type representing some syntactic element, the file range (source file and offset range) from which the element was parsed, and zero or more sub-element children. The AST is unified in that shared language features are mapped to the same representation, as illustrated by Figure 5. Although tree-sitter
does not guarantee perfect parsing, this is tolerable under our approximate design. Our approach also avoids heavyweight and specialized compiler frameworks necessary for precise parsing (e.g., clang
(Lattner and Adve, 2004) for C
, javac
(Arnold et al., 2005) for Java, CPython (van Rossum, 1995) for Python, etc.), significantly simplifying the implementation.
CFG Construction. The next step lowers the AST into a Control Flow Graph (CFG), where each node represents a statement or conditional, followed by one (or more) control-flow edges to successor nodes. The CFG is language agnostic, and common language features (e.g., if-then-else, while-loops, etc.) are lowered into common CFG patterns.
Slice generation.
The path partitioning is generated using Algorithm 3.
For each partition , a corresponding truncated sub-program is generated by substitution non-covered nodes/sub-programs with an assume(0)
statement, followed by simplification. From this, a truncated slice is generated by applying Weiser’s back-slicing algorithm (as illustrated in Algorithm 2). Here, the variables from the post-condition () are used as the slicing criterion, and Algorithm 2 removes all nodes from of which is not control- or data-flow dependent.
4.2. Truncated Slice Rendering
So far, each slice () is represented as a sub-graph of the (truncated) CFG. Our ultimate target is an off-the-shelf LLM, meaning that the slice must be rendered back into a generic text form. For this we leverage the underlying AST representation based on file ranges. For the rendering algorithm, we use an interval tree () mapping ranges to strings. For each CFG node (and corresponding AST statement ), we insert all into .
Context.
We also include the AST path from the root down to each sliced (), such as the enclosing function, control-flow context (if-then-else, while-loops, etc.), or class definitions (for Java). This ensures the rendered slice is coherent (re-parsable) code, rather than a collection of isolated statements. Furthermore, it is often useful to include all relevant non-executable AST nodes from which each depends, including: variable declarations, type declarations, member declarations, function declarations, global declarations, macro definitions, etc. Such dependencies are matched based on the AST name. For example, given the executable statement (xs = insert(xs, x)
), then any function declaration or macro definition matching the name “insert
” can be included in the context. Since we rely on lightweight parsing and not a compiler front-end with semantic information, name-based matching may over-approximate dependencies.
However, this is allowable under our approximate workflow, and the LLM will often ignore irrelevant context.
The slice renderer also preserves the original formatting, including the preceding and succeeding whitespace and comments. Preserving such information is not strictly necessary for C
and Java, but is essential for whitespace-sensitive languages such as Python. Furthermore, source-code comments can provide additional context, such as programmer intent, etc., which can assist LLM inference.
Example.
An example rendered slice is illustrated in Figure 4 (d). Here, all sliced executable statements are preserved, as well as the enclosing context (e.g., function declaration for f()
). Furthermore, the inner if-then-else has been simplified to an assertion, and the formatting (whitespace) has been preserved. The resulting slice is coherent C
code in its own right. Crucially, each path through the slice corresponds to an equivalent path in the original programming, meaning that any inference on the slice is also a valid inference for the original code.
LLM Inference.
Once the rendered slice has been generated, the final step is prompt construction and LLM inference. This step can be highly customized, but for our basic design, we use a prompt structure that mirrors a Hoare triple where is a pre-condition, is a post-condition, and is the generated slice (see Example 2.3, Example 3.2, and Figure 4 (e)). In addition, some basic instructions for the LLM are provided, such as the output format. Here, we assume that the LLM will generate one of two possible responses to the prompt; namely PASS
(the post-condition holds) or FAIL
(the post-condition does not hold)
4.3. Implementation
We have implemented our workflow in the form of the AutoExe tool. AutoExe takes as input a program in a supported programming language (currently C
/Java/Python), pre- and post-conditions ( and ) expressed as code, constraints, or natural language. AutoExe automatically decomposes the program into a sequence of truncated slices, and then constructs a sequence of prompts (a.k.a. verification conditions) to be sent to the LLM for inference. The slices are ordered by size, as to find the least counter-example to the post-condition where applicable. AutoExe is also language agnostic modulo the parser and some elements of the renderer. Furthermore, AutoExe is lightweight and approximate by design—significantly simplifying the implementation (i.e., does not rely on language-specific compiler infrastructure). This also reflects the nature of LLM inference, which is heuristical by nature, rather than relying on precise parsing and semantic analysis, and yet can still make useful inference for many real-world applications.
5. Evaluation
To evaluate the effectiveness of LLM-based symbolic execution implemented by AutoExe, we consider the following research questions:
- RQ1:
Accuracy: What is the accuracy of AutoExe compared to existing methods?
- RQ2:
Scale: Can AutoExe scale to large code bases?
- RQ3:
Language Agnostic: Can the AutoExe workflow support multiple programming languages?
5.1. Experiments Dataset and Baseline
The dataset used in the experiments is gathered from two primary sources. The first source is the REval database (Chen et al., 2024b), which consists of curated Python solutions to LeetCode problems with accompanying test suites. These subjects were selected based on the presence of multiple control flow constructs, including nested conditionals and loops, ensuring a diverse set of program behaviors that cannot be solved easily by traditional symbolic execution methods nor whole-program LLM prompting. Additionally, we curated a dataset of C
-language solutions to LeetCode problems to complement the REval dataset and assess the tool’s ability to generalize across programming languages. These C
solutions were selected following similar criteria, emphasizing functions with multiple control flow branches to ensure the dataset’s representativeness.
One challenge was the construction of non-trivial Hoare triples for evaluation since REval only provides the code and a test suite, while the LeetCode database we used only consists of code and problem descriptions. For this, we use two distinct methods:
Python-Desc: Utilizing the availability of program descriptions, this case used the description of the problem as the natural-language post-conditions of the Hoare triple and also used the restrictions on input (if any) as the pre-condition.
Mixed-Curated: A set of 31 subjects that consists of both Python and C programs curated from both the problem solutions in the original REval dataset and the LeetCode dataset, that comes with handwritten and nontrivial pre- and post-conditions that reflect hidden properties of the original problem and implementation.
We consider two main baselines for comparison. The first is traditional symbolic execution tools. For Python, we considered CrossHair (Schanely, 2017), PyExZ3 (BALL et al., 2015), and pySym (Bann, 2016). However, with the exception of CrossHair, most traditional symbolic execution tools are unmaintained (6 years or more, as of writing). For C
, we consider KLEE (Cadar et al., 2008) as a mature traditional symbolic execution tool. One practical issue with traditional tools is the lack of expressivity in the pre- and post-conditions, which must be specified in either code or a formal language. For some subjects (Python-Desc), the pre-/post-conditions are expressed in natural language, which can be handled by LLMs but not traditional tools.
For the second baseline, we consider an “ad hoc” LLM-based program analysis over the entire subject. For this baseline, we query the LLM for a counter-example (if one exists) or whether the post-condition always holds for inputs that satisfy the pre-condition. While such an approach may yield reasonable results for more straightforward programs, it lacks the systematic partitioning capabilities that are provided by our framework.
Test Models. We tested AutoExe on several open-source LLMs:
Meta’s Llama3 series (Grattafiori et al., 2024): Instruction-tuned models that are fine-tuned and optimized for general use cases like dialogue and chatting.
Microsoft’s Phi-4 (Abdin et al., 2024): A 14B model that is built on high-quality data from filtered public domain websites and acquired academic books, designed to aid research on language models.
DeepSeek’s DeepSeek-R1: A first-generation reasoning model coming from DeepSeek, achieving performance comparable to OpenAI-o1 (DeepSeek-AI et al., 2025) across math, code and reasoning tasks with much smaller parameters.
Google’s Gemma3 (Team et al., 2024): A lightweight family of multimodal models that features a larger context window of 128K.
Note that AutoExe is optimized for local models, since it potentially generates a large amount of prompt traffic. All experiments are run using ollama
version v0.5.7.
5.2. RQ1: Accuracy
To evaluate the accuracy of AutoExe in generating counter-examples and verifying post-conditions, we compare the results of the baseline approach and the output of AutoExe for the above dataset. Specifically, for each program, we executed AutoExe and the baseline approach, recording their respective outputs (verified, unknown, or unverified with a counter-example) under the same prompt and LLM model. Each test is run 5 times, and the most common output is selected as the final result to mitigate random variations in the results. The accuracy of each test is measured by comparing the result against ground truth derived from the provided test suites and manual verification.
Model | Method | Python-Desc | Mixed-Curated | ||||
Total | Correct | Accuracy | Total | Correct | Accuracy | ||
Llama3-8B | AutoExe | 85 | 82 | 96.5% | 31 | 23 | 74.2% |
Baseline | 85 | 79 | 92.9% | 31 | 22 | 71.0% | |
Llama3.1-8B | AutoExe | 85 | 84 | 98.8% | 31 | 24 | 77.4% |
Baseline | 85 | 71 | 83.5% | 31 | 19 | 61.3% | |
Llama3.3-70B | AutoExe | 85 | 72 | 84.7% | 31 | 27 | 87.1% |
Baseline | 85 | 70 | 82.4% | 31 | 22 | 71.0% | |
Phi-4-14B | AutoExe | 85 | 74 | 87.1% | 31 | 24 | 77.4% |
Baseline | 85 | 72 | 84.7% | 31 | 24 | 77.4% | |
DeepSeek-R1-70B | AutoExe | 85 | 85 | 100.0% | 31 | 19 | 61.3% |
Baseline | 85 | 82 | 96.5% | 31 | 17 | 54.8% | |
Gemma3-4B | AutoExe | 85 | 73 | 85.9% | 31 | 20 | 64.5% |
Baseline | 85 | 64 | 75.3% | 31 | 16 | 51.6% | |
Gemma3-27B | AutoExe | 85 | 72 | 84.7% | 31 | 20 | 64.5% |
Baseline | 85 | 76 | 89.4% | 31 | 23 | 74.2% | |
Average | AutoExe | 85 | 77.4 | 91.1% | 31 | 22.4 | 72.4% |
Baseline | 85 | 73.4 | 86.4% | 31 | 20.4 | 65.9% |
Main result. Figure 6 shows the results of evaluating AutoExe on different datasets presented above. Compared against the baseline available, the experimental results demonstrate that AutoExe generally improves accuracy compared to the baseline across the models and datasets being tested, though the extent of improvement varies. For the Python-Desc dataset, AutoExe consistently outperforms the baseline for most subjects and models. Meanwhile, on the Mixed-Curated dataset, improvement in accuracy is less obvious due to the dataset consisting of more complex programs with a lower-level language, but AutoExe still presents similar performance compared to baseline methods, even in such settings. We also note that the baseline accuracy for LeetCode-derived datasets is already high due to the relatively small size of the subjects. We examine some more realistic case studies in RQ2, where the impact of partitioning is stronger.
We also attempted to run CrossHair (Schanely, 2017) on the Python portion of the Mixed-Curated dataset since the post-conditions are expressed in executable Python code. However, CrossHair was unable to identify any counterexample for all Python subjects—likely due to the limitations of traditional symbolic execution as well as the specific implementation. The other subjects, including Python-Desc, express the post-condition as natural language descriptions and thus are unsuitable for CrossHair (or other surveyed traditional symbolic execution tools).
Impact of different LLMs. Figure 6 above also evaluates AutoExe on multiple different LLMs, including Meta’s Llama3, Llama3.1, and Llama3.3 models (Grattafiori et al., 2024) of various parameter sizes, Microsoft’s Phi-4 model (Abdin et al., 2024), 4B and 27B versions of Google’s Gemma3 model (Team et al., 2024), and the 70B distilled version of DeepSeek-R1 model (DeepSeek-AI et al., 2025). Interestingly, the results reveal some notable differences in performance across various LLMs, highlighting the impact of model size and specialization on the results. Among the tested models, the most significant improvement in accuracy is observed with Llama3.1-8B, where AutoExe achieves 98.8% accuracy compared to the baseline’s 83.5%, highlighting the benefits of partitioning in improving the accuracy of LLM reasoning for comparatively small models. In contrast, on larger models like Llama3.3-70B and DeepSeek-R1-70B, such improvements in accuracy are more subtle, mainly due to the fact that larger models inherently have a stronger reasoning ability. This is a positive result—smaller LLMs can be run locally on consumer hardware, and AutoExe exhibits a clear benefit for these use cases. For larger models, the result is still positive, with the exception of Gemma3-27B. We expect that the relatively small size of the test subjects benefits ad hoc analysis. We shall test larger subjects in RQ2.
Method | Python-Desc | Mixed-Curated |
AutoExe Min-Counter TCs | ||
Baseline TCs |
Prompt size.
The statistics for the average Token Counts (TCs) across different datasets are shown in Figure 7. The TCs for each query to the LLM are calculated using tiktoken
(Jain, 2025), with the corresponding Byte-Pair Encoding (BDE) token encoding (also used by GPT-4o mini (OpenAI, 2024)). In the table, we compare the average TCs for the baseline method against the average TCs for the minimum counterexample for each test case in AutoExe’s execution using Llama3.1-8B as the LLM.
From the table, AutoExe demonstrates practical token count reduction ability compared to the baseline method—reducing the length and complexity of the prompt—with reductions of approximately 9.5% and 22.9% on average, respectively, for Python-Desc and Mixed-Curated datasets. We shall also test larger subjects in RQ2.
Case study.
We present Figure 8 as a concrete example. Here, the pre- and post-conditions of the original Python function (shown in Figure 8 (a)) are indicated with the corresponding PRE
and POST
comments. In this example, the original function has a pre-condition that value is not empty, which is indicated by the (assume len(value) > 0
) statement located at the beginning of the function. The post-condition of the function is listed as , which does not always hold, as the given function will round up any decimal input ending in “.5
”. For example, , thus violating the post-condition. We find that most small LLMs cannot correctly analyze the whole program. A typical response of GPT-4o mini (OpenAI, 2024) when fed with the entirety of the original program is:
“Rounding using or will always produce a result that is at least less than or equal to the absolute value of value.”
The LLM then proceeds to incorrectly declare that the post-condition of the program will always be satisfied. In this instance, we speculate that the LLM is reasoning over what the program is supposed to do rather than what the program actually does—i.e., a form of hallucination.
(a) original Python function
(b) truncated slice #1
(c) truncated slice #2
In contrast, AutoExe decomposes the input program into truncated slices. Two of the generated slices of the original Python function are shown in Figure 8 (b) and (c). A key observation of those generated slices are their reduced complexity. For example, in (b), only the elif
branch is considered, with the others truncated, forming implicit pre-conditions in the form of assume
-statements. In (c), only the second if
branch is assumed to be taken. In this example, we can see that AutoExe significantly reduces the complexity and size of the resulting truncated slices. Specifically, the original Python code (a) contains 430 tokens, which is reduced to 103 tokens ( reduction) for (b) and 69 tokens ( reduction) for (c). Together with the reduced CFG complexity, the slices form a smaller and more targeted prompt, increasing the chance of a correct validation result. With the truncated slices, the LLM correctly infers the post-condition holds for (b) but does not hold for (c). For example, GPT-4o mini outputs the following for (c) (emphasis original):
“There are cases (in fact, all cases where
num > 0
and ends in’.5’
) that violate thePOST
condition sinceabs(res)
will always be greater thanabs(float(value))
. Thus, the answer is No, the post-condition (POST
) does not always hold.”
Since there exists a counter-example (c), we have shown that the post-condition does not hold for the original program (a).
We also attempted to analyze Figure 8 (a) using CrossHair (Schanely, 2017), a traditional symbolic execution engine for Python. However, CrossHair terminates after failing to find any violation of the post-condition. We believe this is due to an incomplete search or incomplete solving of the path constraint by the underlying SMT solver (z3). In contrast, AutoExe can exhaustively enumerate all path partitions, and the underlying LLM makes the correct inferences on the corresponding slice. Even this simple case study highlights the challenges with traditional methods.
5.3. RQ2: Scale
Token Count | Llama3.1 | Gemma3 | ||||||||
Subject | Bug | File | Func | Slice | File | Func | Slice | File | Func | Slice |
xinput | NULL-pointer | 4281 | 1197 | 178 | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ |
xlsclients | NULL-pointer | 140120 | 1090 | 229 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ |
xmodmap | NULL-pointer | 18683 | 2189 | 279 | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ |
xset | Divide-by-zero | 17078 | 3164 | 506 | ✗ | ✓ | ✓ | ✗ | ✓ | ✓ |
xwininfo | Buffer-overflow | 152020 | 2560 | 597 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ |
Average | 66436 | 3.07% | 0.54% | 0% | 20% | 60% | 0% | 20% | 60% |
To evaluate the ability of AutoExe to scale, we consider a large read-world project, specifically the X11 (X.Org Foundation, 2024) client libraries and applications, of which several recent crashes were found via fuzzing (Meng et al., 2024b). For these tests, we construct post-conditions by negating the crash condition, e.g., () for NULL-pointer dereference. The test suite consists of K sLOC of C
code (M tokens)—far too large for analysis within a single prompt—meaning that some kind of decomposition is necessary. We consider the following decompositions: (File) only relevant source files included, (Func) only relevant functions included, (Slice) the AutoExe minimum counter-example assessed against the ground truth. The results are shown in Figure 9. We consider five X11 client applications (xinput
, xlsclients
, xmodmap
, xset
, xwininfo
The results are shown in Figure 9. Unsurprisingly, the slice-based decomposition (Slice) is the most effective at reducing the total token count over the naïve (File) and (Func) based methods. The reduction in the prompt size and complexity also has a noticeable impact on LLM inference. For this test, we consider two small models (Llama3.1-8B and Gemma3-4B) and see that (Slice) achieves 60% accuracy (3/5), compared to 20% for (Func) and 0% for (File). We also note that, for larger subjects, the reduction in token counts and accuracy are far more impactful than the relatively small subjects of RQ1. An example slice for xinput
is shown in Figure 10. Once the program has been reduced, the bug becomes apparent: the XIQueryDevice()
function returns NULL
(derived from an error-handling path), and this value is immediately passed to XIFreeDeviceInfo()
(the intermediate code was removed by the slicer), thus violating the post-condition. After the truncation and slicing, the violation is readily apparent to even small LLMs. In the non-sliced code, the bug is significantly more obfuscated by additional (and irrelevant) code. For the original code, see here and here.
5.4. RQ3: Language Agnosticism
AutoExe implements a lightweight workflow that is language agnostic. To evaluate AutoExe’s ability to analyze different programming languages, we analyze the same subjects implemented in Python, C
, and Java. For this, we use GPT-4o-mini to automatically translate the Python-Desc dataset into equivalent C
and Java programs. The translated program is then gathered to form the Translated-C-Desc and Translated-Java-Desc datasets, also consisting of 85 programs. Note that the translation is not necessarily faithful, but can still be processed by AutoExe using the same evaluation framework detailed in RQ1. The results are summarized in Figure 11.
Model | Method | Translated-C-Desc | Translated-Java-Desc | ||||
Total | Correct | Accuracy | Total | Correct | Accuracy | ||
Average | AutoExe | 85 | 73.3 | 86.2% | 85 | 76.0 | 89.4% |
Baseline | 85 | 66.5 | 78.2% | 85 | 71.3 | 83.8% |
Compared against the original Python dataset, we can see there is a general decline in accuracy for both baseline methods and AutoExe. Nevertheless, AutoExe is still able to improve accuracy compared to the baseline method for the translated datasets, demonstrating its versatility in handling multi-language programs that is the direct result of its language-agnostic workflow and the inherent ability for LLM to understand multiple languages. These results are achieved with reliance on precise parsers and/or language-specific compiler front-ends.
5.5. Discussion
Our results show that path-based decomposition of program analysis tasks is effective at improving LLM-based program analysis, especially for small LLMs and real-world problems. The significance is that it allows for higher accuracy to be achieved with smaller models that can be run on consumer-grade hardware. In addition, we prove that a lightweight and language agnostic workflow is feasible and can still achieve good results.
Limitations. LLM-based program analysis is applicable to program analysis tasks that cannot be handled by traditional means. That said, the probabilistic reasoning on LLMs is not suitable for all applications, even with improved accuracy. LLMs are also unlikely to ever achieve the same accuracy as traditional solvers for some tasks, such as solving systems of linear equations. As such, we propose LLM-based symbolic execution as a complementary method that does not necessarily replace traditional approaches for all use cases. Another limitation is that path-based decomposition may still explode, even when our approach is guaranteed to terminate. This is an inherent limitation of path-based reasoning. However, we believe that the truncated-slice based decomposition is a significant mitigation.
6. Related Work
Static program analysis via symbolic execution.
Symbolic execution (King, 1976) is an established method for static programs whose origins also relate to early work formalizing programs as mathematical logic. The idea is to execute symbolic states, representing sets of concrete inputs, allowing for the exhaustive exploration of program behavior. Over the decades, many different symbolic execution engines and frameworks have been developed, such as: KLEE (Cadar et al., 2008), Owi (Andrès et al., 2024), Symbolic PathFinder (SPF) (Păsăreanu and Rungta, 2010), Java PathFinder (JPF) (Visser et al., 2004), CrossHair (Schanely, 2017), Angr (Shoshitaishvili et al., 2016), S2E (Chipounov et al., 2012), PyExZ3 (BALL et al., 2015), etc. Unlike our approach, such traditional tools translate paths into some underlying formal language for theorem proving and thus inherit many of the limitations discussed in this paper. Furthermore, most existing tools are specialized to a specific language (C
, Java, binary, etc.) and are closely integrated into specific compiler frameworks (e.g., LLVM (Lattner and Adve, 2004) for KLEE). That said, LLMs use a fundamentally different type of reasoning compared to the deductive reasoning of theorem provers. As such, traditional approaches are suitable for problems that can be handled by traditional methods and for applications where perfect accuracy is required.
LLM-based program analysis. One recent alternative to traditional program analysis methods is Large Language Models (LLMs). LLMs are very general tools and can be applied to a wide variety of tasks, including fuzzing (Asmita et al., 2024; Meng et al., 2024c), vulnerability detection (Zhou et al., 2024b), and program repair (Fan et al., 2023). Another recent innovation is LLM-based agents (Xi et al., 2025), which are algorithms where decisions are made by the LLM. Our core Algorithm 1 is traditional and not agent-based. However, an agentized version could be made, but decisions (e.g., which branch to explore first) could be deferred to the LLM.
Intersection between symbolic execution and LLMs. There are some other nascent works combining LLMs and symbolic execution. LLM-Sym (Wang et al., 2024) is an agent-based symbolic execution framework for Python code that uses an LLM to translate paths into traditional path constraints suitable for solving via z3 (de Moura and Bjørner, 2008). LLM-Sym is fundamentally different in that our approach avoids translation altogether, instead directly using the LLM itself as a solver. Since LLM-Sym still uses translation to z3, it still inherits many of the limitations of traditional symbolic execution engines discussed in this paper. Similarly, Loopy (Kamath et al., 2023) aims to discover loop invariants using LLMs, which can then be applied to symbolic analysis. Our approach avoids the need for invariant discovery since it is not based on translation.
7. Conclusion
Large Language Model (LLM)-based program analysis has enjoyed an explosion of application over the past few years. The ability of LLMs to reason over code directly has advantages over traditional static program analysis methods, such as symbolic execution, that have significant limitations regarding difficult-to-analyze programs (e.g., loops, environment, heaps, etc.). That said, LLMs also have inherent limitations. For example, since LLMs essentially use a form of approximate or probabilistic reasoning, they are not always accurate, which may limit their application to certain tasks. Another limitation is scale, where the LLM token limit restricts the size and scope of the program to be analyzed. Such problems can be mitigated using a larger LLM, but this has additional costs, such as the requirement of enterprise grade GPUs in order to run the model.
In this paper, we introduce a variant of symbolic execution that uses an LLM as the underlying reasoning engine instead of a traditional theorem prover or SMT solver. Our approach introduces a generic path constraint representation in terms of the original code—allowing the LLM to reason directly over the path constraint and avoiding translation into a (less expressive) formal language. Our approach allows for a path-based decomposition of the analysis task into smaller (more tractable) sub-tasks, which use less tokens (helping scale), and are more targeted (helping accuracy). We implemented our approach in the form of AutoExe—a practical LLM-based symbolic execution engine that supports multiple programming languages (i.e., language agnostic, supporting C
/Python/Java) without depending on heavyweight compiler infrastructure. Our experimental results demonstrate measurable improvements in terms of both accuracy and scale, especially in smaller models that can run on consumer grade GPUs.