### v2

# Large Language Model Powered Symbolic Execution

Abstract.
Large Language Models (LLMs) have emerged as a promising alternative to traditional static program analysis methods, such as symbolic execution, offering the ability to reason over code directly without relying on theorem provers or SMT solvers. However, LLMs are also inherently approximate by nature, and therefore face significant challenges in relation to the accuracy and scale of analysis in real-world applications. Such issues often necessitate the use of larger LLMs with higher token limits, but this requires enterprise-grade hardware (GPUs) and thus limits accessibility for many users. In this paper, we propose LLM-based symbolic execution—a novel approach that enhances LLM inference via a path-based decomposition of the program analysis tasks into smaller (more tractable) subtasks. The core idea is to generalize path constraints using a generic code-based representation that the LLM can directly reason over, and without translation into another (less-expressive) formal language. We implement our approach in the form of AutoBug, an LLM-based symbolic execution engine that is lightweight and language-agnostic, making it a practical tool for analyzing code that is challenging for traditional approaches. We show that AutoBug can improve both the accuracy and scale of LLM-based program analysis, especially for smaller LLMs that can run on consumer-grade hardware.
1. Introduction
Program analysis is a foundational discipline in computer science that aims to understand program behavior through various systematic techniques. Traditional forms of program analysis include static methods—such as symbolic execution (King, 1976), abstract interpretation (Cousot and Cousot, 1977, 1979), and model checking (Clarke and Emerson, 1982)—that analyze the program without executing it, as well as dynamic methods—such as fuzzing (Miller et al., 1990; Zeller et al., 2024), concolic execution (Godefroid et al., 2005; Sen et al., 2005), instrumentation (Serebryany et al., 2012; Nethercote and Seward, 2007), and profiling (Graham et al., 1982)—that analyze program behavior based on observations of actual executions. In general, program analysis has many applications, such as program testing (Hamlet, 1977; Clarke, 1976; Rapps and Weyuker, 1985; Zhu et al., 1997; Godefroid et al., 2005; Cadar et al., 2008; Sen et al., 2005; Pacheco and Ernst, 2007; Weyuker, 1983; Goodenough and Gerhart, 1975), debugging (Zeller, 2005; Agrawal et al., 1993; Engler et al., 2001), verification (Floyd, 1967; Hoare, 1969), repair (Le Goues et al., 2012; Long and Rinard, 2016), reverse engineering (Nelson, 2005), and vulnerability detection (Kaur and Nayyar, 2020). A recent alternative to traditional program analysis methods has emerged in the form of Large Language Models (LLMs) (Chen and et al., 2021; Li et al., 2022; Touvron et al., 2023; Bubeck et al., 2023; OpenAI, 2024). Here, LLMs can make inferences over code directly with properties expressed in code or natural language and have become powerful enough to handle many traditional program analysis applications. As such, many LLM-based testing (Chen et al., 2024a; Meng et al., 2024c), debugging (Majdoub and Charrada, 2024), and repair (Zhou et al., 2024a, b; Xia et al., 2023) tools have recently emerged.
We contrast traditional program analysis (symbolic execution) with LLM-based program analysis. Here, symbolic execution is a static program analysis method based on the idea of executing a program with symbolic inputs/values, typically represented as logical formulae over some underlying theories (e.g., linear arithmetic, bit vectors, and arrays). Symbolic execution systematically transforms these symbolic values (a.k.a. the symbolic state) based on the program’s statements, effectively executing multiple concrete paths simultaneously. These symbolic states can then be used for various analysis tasks, such as verifying the truth of an assertion or generating test cases, using deductive inference with the help of some underlying theorem prover or SMT solver (de Moura and Bjørner, 2008). In contrast, LLM-based program analysis involves engineering a prompt (or sequence of prompts) that queries an LLM as an oracle. A typical prompt consists of the relevant code (or code fragments), as well as instructions explaining the analysis task expressed in natural language. The LLM effectively uses a form of approximate inference (based on the training data) to solve the task, rather than the strict deductive inference used by a theorem prover. As such, LLM-based program analysis is typically ad hoc and tailored to specific tasks rather than general and principled.
Both traditional and LLM-based program analysis face significant challenges in practice. For example, traditional symbolic execution has several well-known limitations (Baldoni et al., 2018; Cadar and Sen, 2013), such as the handling of unbounded loops, the handling of external environment/libraries, and the handling of memory/heap-manipulating programs—all of which are common-place in real-world code. We consider how KLEE (Cadar et al., 2008), a prominent symbolic execution engine, treats each iteration of a loop as a separate path, leading to non-termination for unbounded loops (e.g., ). In contrast, LLMs have the ability to directly reason over loops, environment, and heaps, avoiding problems such as non-termination. That said, LLMs face other challenges, such as scalability (e.g., the 8192 token limit for GPT-4 (Bubeck et al., 2023)) and accuracy (Levy et al., 2024; Fang et al., 2024) due to approximate reasoning. Recent studies also correlate the accuracy of LLMs with prompt size (Levy et al., 2024), meaning that more concise and targeted prompts generally perform better.
In this paper, our aim is to improve the accuracy and scale of LLM-based program analysis. Our first main insight is that the strengths/weaknesses of traditional and LLM-based program analysis are complementary, meaning that a hybrid design can help address the limitations of either approach. To this end, we propose combining the path-based decomposition of symbolic execution with approximate inference via an LLM—a.k.a. LLM-based symbolic execution. The core idea is a principled decomposition of the original program analysis task into smaller subtasks based on paths in the original program—helping to mitigate some of the scalability and accuracy concerns with LLMs. Our second main insight is that since LLMs are primarily trained on code, the representation of symbolic states should also be in terms of code rather than logical formulae. To this end, we propose a generic path constraint representation of the form of a strongest post-condition () predicate transformer (Dijkstra, 1975) over sub-programs derived from the original program, where each sub-program represents a path. Since we represent path constraints as ordinary code, we can use an LLM prompting directly, disposing of any verification conditions (VCs) generated by the symbolic execution process. Our approach also avoids many of the limitations associated with the translation of paths into formulae for a theorem prover (e.g., environment and heaps). Our final insight is that our path constraint representation can be generalized into sets of paths (e.g., all iterations of an unbounded loop), ensuring that LLM-based symbolic execution will always terminate.
We study LLM-based symbolic execution in both theory and practice. We show that LLM-based symbolic execution mitigates many of the limitations of both traditional and LLM-based program analysis. We have implemented our approach in the form of AutoBug—an automated LLM-based symbolic execution engine. AutoBug uses a path-based decomposition of a program analysis task into smaller (more tractable) subtasks that are suitable for LLM inference. Based on the observation that LLMs are inherently approximate oracles, we propose a lightweight design of AutoBug that is language agnostic and does not rely on any heavyweight compiler infrastructure. AutoBug is designed to be a practical program analysis tool that can be applied to code that is difficult to analyze with traditional methods. In summary, the main contributions of this paper are:
We introduce the concept of LLM-based symbolic execution—a symbolic execution methodology using LLMs for direct reasoning over the original programming language (e.g.,
C
/Java/Python), rather than indirect reasoning via translation into some formal theorem prover language. We show that our approach avoids many well-known limitations of traditional symbolic execution, such as unbounded loops, external environment, heap manipulation, etc. Furthermore, our approach uses a path-based decomposition of program analysis tasks into more concise LLM prompts, improving the accuracy and scalability of LLM inference. -
We implement our approach in the form of AutoBug—a lightweight LLM-based symbolic execution engine that supports multiple programming languages without relying on heavyweight compiler infrastructure.
We evaluate AutoBug against various program analysis tasks for
C
/Python/Java code. We show that AutoBug improves both accuracy and scale, especially for smaller models that can run on consumer-grade hardware.
2. Motivation
2.1. Background
2.1.1. Symbolic Execution
Symbolic execution (King, 1976; Baldoni et al., 2018) is an established program analysis methodology based on running the program with symbolic values representing sets of concrete inputs (rather than a single concrete input for normal execution). For each symbolic input, symbolic execution will systematically explore multiple execution paths of the program, allowing for applications such as error detection, verifying correctness, or finding vulnerabilities.
Symbolic execution works with symbolic states
that are traditionally represented as a set of variables (e.g., x
and y
) subject to a path constraint (e.g., ).
The symbolic state represents the set of concrete states (e.g., ) that is reachable from some symbolic input.
Symbolic execution can be defined in terms of operations over symbolic states, e.g., symbolic execution over C
-style increment statement (x++
) can be represented as the Hoare triple (Hoare, 1969):
Given a pre-condition (), the triple describes the resulting post-condition after the execution of the increment operation. For example, given a path constraint , then the path constraint after the operation will be , or equivalently, . Conditional statements (e.g., ) are typically handled by forking the path constraint () into separate and constraints, and then continuing execution along the individual then- and else-branches. A similar strategy is used for loops (e.g., ). Finally, for each path through the program, a final path constraint will be generated. This can be used to prove properties over the path, such as whether a final post-condition () holds, i.e., whether the verification condition (VC) of the form () holds or not. Such VCs are discharged with the help of theorem provers, such as SMT solvers (de Moura and Bjørner, 2008). By executing sets of concrete inputs at once, symbolic execution aims to exhaustively explore the space of program behaviors—something that cannot be achieved by concrete execution alone.
Example 2.0.
Consider the simple program () and the post-condition . Then the following VC will be generated for the -path:
(VC-Else) |
A theorem prover shows that (VC-Else) is unsatisfiable, meaning that () does not hold for all executions of the program. ∎
2.1.2. Large Language Models
The emergence of Large Language Models (LLMs) (Chen and et al., 2021; Li et al., 2022; Touvron et al., 2023; Bubeck et al., 2023; OpenAI, 2024) presents a new alternative for reasoning over program code.
Trained on extensive datasets, including millions of lines of real-world code written in multiple languages, LLMs can reason over C
/Java/Python/etc. code directly, including many traditional program analysis tasks, such as
testing (Chen et al., 2024a; Meng et al., 2024c), debugging (Majdoub and Charrada, 2024), and repair (Zhou et al., 2024a, b; Xia et al., 2023).
For example, instead of relying on symbolic execution, LLMs can reason over code using a suitable prompt expressed in natural language, such as “there is a bug present in the following Python code segment, please suggest the possible root causes of the bug and corresponding fixes”.
Most modern LLMs can analyze the code and generate possible suggestions and patches automatically, based on the understanding of code and defects present in the training set.
Unlike traditional program analysis methods, LLMs do not aim to be perfectly precise. Rather, LLMs can be thought of as approximate oracles that are sometimes incomplete or give the wrong answer. This is because LLMs fundamentally rely on learned patterns and approximate reasoning, rather than classical deductive reasoning used by traditional theorem provers. Despite the difference, LLMs are clearly useful, with an explosion of applications for real-world analysis problems.
(a) original program
(b) slice for “mutx is unlocked”
(c) slice for “”
2.2. Limitations of Symbolic Execution
While symbolic execution has many applications (e.g., bug detection, security analysis, debugging, and program repair), it also has well-known limitations, as we briefly summarize below:
2.2.1. Limitation: Handling Loops (and Recursion)
Unbounded loops (and recursion) are a known problem for symbolic execution. Here, symbolic state forking (Section 2.1.1) can treat each loop iteration as a new path, potentially leading to infinite unrolling if the loop is not bounded. Popular symbolic execution tools, such as KLEE (Cadar et al., 2008), handle this problem using a concrete iteration bound—replacing potentially infinite exploration with a bounded, but incomplete, exploration.
An alternative is to use loop invariants. If known, the invariant allows symbolic execution to pass over a loop without explicit unrolling. Loop invariants can be manually provided, or discovered automatically, such as using abstract interpretation (Cousot and Cousot, 1977) over some known domain, constraint based (CBMC (Gadelha et al., 2017)), or machine learning (Code2Inv (Si et al., 2018)). However, loop invariant discovery may only work for simple loops, and the general case is either computationally hard or undecidable.
2.2.2. Limitation: Handling External Environment
Another known problem is the handling of external functions (e.g., calls to third-party libraries without source code) and/or external inputs (e.g., recv
from a socket), collectively the external environment.
Since the underlying symbolic-execution theorem prover uses deductive reasoning, a precise specification of all external operations and/or inputs is usually required.
As such, the environment is usually handled through a combination of stubs, modeling, or concretization.
For example, the user can manually model an external function call by implementing a replacement stub function that specifies the necessary specification using klee_assume()
However, this approach is manual, and modeling arbitrary code or inputs can require significant effort, meaning that the approach tends to rarely scale.
Another approach is concretization, where the symbolic execution algorithm assigns concrete values to some symbolic variables, allowing external functions to be executed with these values. However, concretization can also lead to an incomplete exploration of program behavior.
2.2.3. Limitation: Heap Manipulating Programs
Traditional theorem provers and SMT solvers tend to have limited support for reasoning over (mutable) data structures with complex structural invariants, such as singly- and doubly-linked lists, binary trees, red-black trees, and DAGs. By extension, traditional symbolic execution tools inherit these limitations. Some tools are based on Separation Logic (Reynolds, 2002), which does support reasoning over heap-manipulating programs, as used by VeriFast (Jacobs et al., 2011) and Infer (Distefano et al., 2019). However, VeriFast is manual and annotation-heavy, and Infer uses heuristics to infer common structural invariants, such as linked lists, but not arbitrary heap shapes.
Discussion
Figure 1 (a) is an example of a program that exhibits all three limitations, including unbounded loops (while (true) ...
), interaction with the external environment (recv
/send
), and is a heap manipulating program (NODE
Although this program is relatively simple, it still presents a significant challenge for traditional symbolic execution tools such as KLEE.
2.3. Limitations of LLMs
LLMs are trained on a huge corpus of data, and do not necessarily have the same limitations as symbolic execution. For example, an LLM can reason over Figure 1 (a), and can answer queries about the program. That said, LLMs also have known limitations, as summarized below:
2.3.1. Limitation: Scale
LLMs typically have a limited ability to reason over large and complex code bases. For example, the token limit imposes a maximum number of tokens for the given input (and output), and a medium-to-large code base can easily exceed this limit.
2.3.2. Limitation: Approximate Oracles
LLMs rely on approximate reasoning, i.e., answers based on statistical patterns and heuristic generalization, rather than the formal deductive reasoning of traditional theorem provers. Even if the token limit (2.3.1) is not exceeded, studies have shown that LLMs generally perform worse with overly verbose prompts that include irrelevant information (Levy et al., 2024).
Discussion
Although scale and accuracy are concerns, LLMs can typically handle more classes of programs than traditional analysis methods, such as symbolic execution. Furthermore, the relative completeness of LLMs, in the absence of precise specifications, means that LLMs can be easily applied to a wide range of applications. The completeness/applicability can often be of greater pragmatic interest than perfect accuracy, and this is one reason behind the explosion of real-world applications.
2.4. Our Approach
Many limitations of traditional symbolic execution stem from those of the underlying theorem prover.
Specifically, existing theorem provers and SMT solvers only accept queries in some formal input language, with limited expressiveness compared to the original source language (e.g., C
/Java/Python).
For example, an SMT solver will only accept queries in the form of a Boolean formula over a given set of theories (T), such as linear inequalities, bit vectors, or arrays.
In contrast, the C
programming language is significantly more expressive, with complex control-flow (loops), memory, pointers, library calls, data structures, environmental interactions, etc.
The discrepancy in the expressiveness complicates the translation from the high-level programming language into the solver input language—such as “unrolling” loops into flat, quantifier-free formulas—and such translation may also be incomplete (e.g., unbounded loops).
Furthermore, the SMT solver may not support the necessary theories for reasoning over complex programs, such as heap-manipulating data structures and external environment interactions.
Our underlying approach is to use the path-based decomposition of symbolic execution, but to replace the traditional theorem prover with an LLM. The key advantage of LLMs is that they can reason over the source code directly—eliminating the need for translation into a less expressive solver input language and the associated limitations. Instead, our approach represents the path constraint as a generic strongest post-condition predicate transformer (Dijkstra, 1975) over a pre-condition , and a derived sub-program which represents a path or set of paths. Our key insight is that LLMs can reason over -constraints directly, since is just ordinary source code, without the need for “translation”. Essentially, we can view the LLM as an effective solver for untranslated -constraints “as-is”. We demonstrate this concept with a simple example.
Example 2.0 (Simple).
Consider the simple program () and post-condition once more (see Example 2.1). Then the path constraint for the -branch can be represented as a formula () or a sub-program (), as follows:
Formula () | Strongest Post-Condition over a Sub-program () |
[…]
The key idea is that both () and are equivalent representations of the same path constraint—i.e., one can be derived from the other under the definition of the language semantics (). While the formula is suitable for a theorem prover, the sub-program is suitable for an LLM:
(VC-Else-2) | “Given , does the post-condition always hold?” |
The LLM determines the VC does not hold, e.g., the following GPT4 (Bubeck et al., 2023) output (emphasis original):
“The post-condition does not always hold. A simple counterexample is when , where instead of . Hence, the claim is false.”
Our approach is to enumerate all sub-programs based on a partitioning of paths through the original program. Here, each sub-program is algorithmically derived from the original code, and contains all statements that (1) are visited by any path from the partition, and (2) of which post-condition is data- or control-flow dependent. Like traditional symbolic execution, our approach is a path-based decomposition of the original program analysis problem into smaller (more tractable) sub-problems. This decomposition helps to address some of the limitations of direct LLM-based reasoning (Section 2.3), such as scale and accuracy. Furthermore, since our approach uses an LLM, it avoids many of the limitations of traditional symbolic execution (Section 2.2) caused by translation. We summarize the benefits as follows:
Handling Loops (and Recursion). Our approach avoids translation of (unbounded) into a less expressive language, and loops/recursion can be represented “as-is” in the derived sub-program(s). Similarly, our approach does not need explicit loop invariant recovery or annotation, as LLMs are capable of reasoning over loops without any special intervention.
Handling External Environment. Rather than manual modeling or concretization, our approach is to use the LLM to infer the likely behavior of the environment or external function call. Since LLMs are trained on a huge corpus of real-world code, they have significant exposure to common libraries, file formats, protocols, etc. Furthermore, even if the external environment is novel, LLMs can still infer the most likely behavior based on clues from the context (function names, variable names, code comments, placement within the algorithm, etc.), as a form of abductive reasoning, or inference to the best explanation, without the need for explicit modeling.
Heap Manipulating Programs. LLMs can directly interpret heap-manipulating programs without the need for any special logical framework. LLMs can also (abductively) infer the (likely) structural invariants based on direct interpretation of the code, without explicit annotation.
Scale. Like traditional symbolic execution, our approach decomposes program analysis problems into smaller (tractable) sub-problems, helping to avoid any hard or soft limit of the LLM. This allows our approach to scale to large/complex programs and analysis problems.
Approximate Oracles. Studies (Levy et al., 2024) show that LLMs perform better with more targeted and concise prompts. By decomposing program analysis problems into sub-problems that capture only the relevant parts of the original (possibly large) code base, we help to focus the LLM and improve the overall accuracy of the analysis.
The decomposition and lack of translation mean that our approach can handle programs that are difficult for traditional program analysis. We illustrate with an example.
Example 2.0 (LLM-based Symbolic Execution).
Consider the Figure 1 (a) program that cannot easily be handled by traditional symbolic execution methods (Section 2.2),
and a simple program analysis problem that verifies each lock(&mutx)
operation is paired with an unlock(&mutx)
operation.
Furthermore, assume that (for the sake of example) the (a) program is too complex for an LLM to handle directly (Section 2.3).111This is not necessarily true, but is an assumption for the sake of an example that can fit within the page limit.
We can express this as a natural language pre- and post-condition ( and respectively) that “mutx
is unlocked”.
Then Figure 1 (b) is an example of a derived sub-program that acts as a substitute for the original program with respect to and .
We have that:
Thus, to refute for (a), it is sufficient to refute for (b).
Furthermore, program (b) is targeted to the specific program analysis task (that mutx
is unlocked), and is smaller in token count (419 vs 81).
The LLM determines the post-condition holds for (b).
Another example is Figure 1 (c) and the post-condition ().
Assuming the same condition initially holds, then (c) is smaller (419 vs 149 tokens).
The LLM determines the post-condition does not hold for (c) since strdup()
may return NULL
Algorithm
Input: A Hoare triple Output: HOLDS or a counterexample 1 2 3Function LLMSymExe(): 4 5 6 7 for do 8 9 10 11 "does hold?" 12 13 if then return 14 15 return HOLDS 16 171ex Algorithm 1 LLM-based Symbolic Execution
An overview of our LLM-based symbolic execution algorithm is summarized in Algorithm 1.
Here, the algorithm’s frontend is similar to that of a standard compiler, and parses the program into an Abstract Syntax Tree (AST, line 2), and then generates a Control Flow Graph (CFG, line 3).
Next, the algorithm generates a representation of the set of all paths through the CFG (line 4).
Here, the set of all paths is represented as a set of path such that (), where each partition represents some (possibly infinite) subset of paths.
One challenge is how to generate a good set of partitioning (to be discussed in Section 3).
Next, for each partition , the algorithm generates a derived sub-program that generalizes the partition (line 6).
For example, Figure 1 (b) and (c) are possible derived sub-programs of Figure 1 (a).
Each sub-program is used to construct a corresponding prompt (line 7), including rendering the back into a text-based source-code representation (line 8).
The prompt (line 7) is a natural language representation of the Hoare triple , which holds iff .
Finally, the prompt is sent to the LLM for inference (line 10).
Algorithm 1 can either establish or refute the post-condition, subject to the reasoning capabilities of the underlying LLM.
Assuming that the partitions are ordered based on slice, Algorithm 1 will return the least sub-program that is deemed to refute the post-condition.
Otherwise, if no such refutation is found, Algorithm 1 deems that the triple holds (HOLDS
Property | Tradition Symbolic Execution | LLM-based Symbolic Execution | |
Design | Overall method | Decomposition & solving path constraints | Decomposition & solving path constraints |
[…]
Specification language | Formal language | Either formal, code, or natural language | |
Cap. | Weaknesses | Loops, environment, heaps | Complex integer, linear, Boolean reasoning |
Unbounded loops | Infinite unfolding or loop invariants | LLM reasons over loops “as-is” | |
Environment | Manual modeling/specifications required | LLM abductive reasoning or training data | |
Heap manipulation | Manual annotation + Separation Logic | LLM reasons over heaps “as-is” | |
Imp. | Programming languages | Language specific (C+KLEE (Cadar et al., 2008) and Java+SPF (Păsăreanu and Rungta, 2010)) | Programming language agnostic |
Compiler infrastructure | Close integration (LLVM (Lattner and Adve, 2004)+KLEE, etc.) | Lightweight (AST-level) implementation |
Summary
Like traditional symbolic execution, LLM-based symbolic execution (Algorithm 1) represents a path-based decomposition the original program analysis task into smaller subtasks. A summary of the main similarities and differences is shown in Table 1. The substitution of a theorem prover with an LLM changes to various aspects of the design, capabilities, and implementation of the symbolic execution engine. For example, LLMs use approximate and abductive reasoning, or rely on information learned during the training process, meaning that LLMs do not need precise specifications or environment modeling. Likewise, the lack of translation into some (less expressive) formal language allows the LLM to reason over loops or heap manipulation “as-is”, without relying on loop/data-structure invariant discovery.
In this paper, we study the concept of program analysis via LLM-based symbolic execution. First, we study the principles of LLM-based symbolic execution in terms of the idealized procedural programming language used by Hoare logic (Hoare, 1969). We show that program analysis tasks can be decomposed into tasks over derived sub-programs representing paths, or sets of paths (truncated slices), through the original program. Furthermore, we also show that only a finite number of (Algorithm 1, line 4) needs to be considered, ensuring that LLM-based symbolic execution will always terminate—even for unbounded loops.
In addition, we study the application of LLM-based symbolic execution in practice.
For this, we design AutoBug—an LLM-based symbolic execution engine for real-world programming languages such as C
/Java/Python.
Our approach is based on the observation that Algorithm 1 is mostly language agnostic except for specific aspects, such as the parser.
This means our approach can be readily ported to other programming languages.
Furthermore, we also observe that, since LLMs are fundamentally approximate, we can build a lightweight implementation that uses approximate parsing and dependency analysis—without relying on any heavyweight and/or language-specific compiler framework.
3. Principles of LLM-Based Symbolic Execution
Our goal is to adapt traditional symbolic execution methods to LLMs that reason over code directly, rather than translation into a (less expressive) theorem prover input language.
3.1. Symbolic Execution Foundations
We use the minimal imperative language defined by Hoare logic (Hoare, 1969) augmented with an explicit -statement.222Note can be emulated as () under partial correctness, but is treated as a special case. We define the language syntax as follows:
Where represents some base language (e.g., arithmetic expressions) and represents Boolean expressions over . We also use to sometimes represent an empty program (equivalent to ). The language semantics are defined inductively (i.e., least relation) in terms of the strongest post-condition () relation defined as follows:
We define a linear program to be any program comprising only -, -, and assignment-statements (without conditionals or loops). We define a to be a linear program that is derived by unfolding the original program using the following rules:
We abstractly define symbolic execution to be any algorithm that combines path unfolding with Verification Condition (VC) solving. Given a program analysis task represented as a Hoare (Hoare, 1969) triple , then a symbolic execution algorithm:
exhaustively generates the set of all through the program; and
solves each corresponding VC () for each .
That is, a symbolic execution algorithm computes the following using explicit enumeration:
The triple is deemed to hold if each individual VC holds for the corresponding path (), and to not hold otherwise. Here, each is defined to be the path condition for the corresponding path . Essentially, a symbolic execution algorithm decomposes the original program analysis task into simpler subtasks that can be solved separately.
Traditional Symbolic Execution
“Traditional” symbolic execution algorithms solve each VC using a suitable theorem prover, such as an SMT solver (de Moura and Bjørner, 2008). To do so, the -constraints for each path are translated into a logical formula over the domain of , by applying the (linear program subset of the) -rules defined above. The translated VC () is then solved using a theorem prover. Note that the translation is necessary since traditional theorem provers are limited to a specific input language (e.g., SMT-LIB), and cannot make inferences over an abstract -constraint directly. In addition, most practical symbolic execution tools implement several optimizations, including incremental path unfolding, incremental -translation, pruning infeasible paths, and the merging of similar paths. For example, rather than pre-computing the set of paths upfront, most practical implementations maintain a symbolic state comprising a current location, a (partially constructed) path constraint/condition, and a set of current variables-of-interest. Such optimizations are consistent with abstract symbolic execution (SymExe) defined above.
We can also understand the limitations of Section 2.2. Firstly, the number of paths in the set may be very large (exponential) or even infinite (unbounded loops, Section 2.2.1). This is known as the path explosion problem, and is a well-known limitation of symbolic execution methods. Secondly, the -translation will be limited for programs that contain operations that are not supported—such as common interactions with the external environment (Section 2.2.2). Finally, the underlying theorem prover itself may be limited. For example, KLEE (Cadar et al., 2008) uses the Z3 (Baldoni et al., 2018) SMT solver over the domain of linear arithmetic, arrays, and bit-vectors by default. However, this configuration does not support reasoning over heap-manipulating programs (Section 2.2.3).
(a) simple program () | (b) path () | (c) truncation () |
1i 2 while i n do 3 4 if x then 5 xs. 6 7 else 8 xs. 9 10 z 11 12 i 13 | 1i 2 3 4 5 xs. 6 z 7 8 i 9 10 11 12 xs. 13 z 14 15 i 16 | 1i 2 while i n do 3 4 5 xs. 6 z 7 8 i 9 (d) slice ( 1i 2 while i n do 3 4 xs. 5 i 6 |
3.2. LLM-Based Symbolic Execution (Path-Based)
The basic idea behind “LLM-based” symbolic execution is to use a Large Language Model (LLM) as the underlying reasoning engine instead of a theorem prover. Thus, given a VC of the form (), we can use the LLM to reason directly over the path constraint and post-condition . This is possible since, through its training process, the LLM can interpret both the syntax of the path (represented as ordinary code), as well as the language semantics represented by the -rules. We illustrate with the following example.
Example 3.0 (Path-based LLM-based Symbolic Execution).
We consider the simple example shown in Figure 2 (a), which is based on Example Program 3 in (Korel and Laski, 1988) . We assume that xs is a data structure implementing a multi-set, and // can be expanded to sub-programs with a suitable multi-set implementation (e.g., a singly linked list), and is initially empty. Furthermore, we assume that the () operation always returns a positive number, which can be expressed in natural language or as a formal rule (). Under these assumptions, the final size of xs should be equal to n. We can express this as the following triple:
Although conceptually simple, the Figure 2 (a) program is challenging for several reasons, including an unbounded loop, environmental input (), and data-structure reasoning (xs). We may prove (Triple) by enumerating paths, such as () from Figure 2 (b), representing two iterations of the loop. We can encode the VC () as a prompt, which is confirmed using an LLM:
“Given the pre-condition and the code , does the post-condition hold?” ∎
This example shows how LLMs can solve tasks that are challenging for traditional program analysis methods. That said, a simple path-based decomposition still inherits some limitations. Firstly, there can still be an infinite number of paths (i.e., path explosion), leading to non-termination. Secondly, each individual path could still be too long for the LLM to effectively reason over (e.g., path from Figure 2 (b) can be generalized to any length). Finally, paths can accumulate irrelevant statements, such as variable z, that can also exacerbate the path length problem.
3.3. LLM-Based Symbolic Execution (Slice-Based)
To address the path-explosion problem, our core idea is to merge individual -constraints (representing individual paths) into generalized -constraints representing (possibly infinite) sets of paths. Given a partition , our approach constructs a truncated sub-program such that:
Thus, instead of disposing of a (possibly infinite) number of verification conditions (VCs) of the form () for each , our approach disposes of a single VC of the form () for the entire set (). First, we shall present a method for constructing a truncated sub-program for a given partition. Next, we shall present a partitioning algorithm that needs only consider a finite number of subsets, even for programs with infinite paths (unbounded loops), ensuring the symbolic execution algorithm always terminates.
3.3.1. Construction
We consider the construction of truncated sub-programs.
Truncation. Given a program and a (possibly infinite) subset of , we derive a truncated sub-program equivalent to for all , and is unreachable otherwise:
Here, () excludes all paths that terminate abnormally via a special (i.e., assume false) statement. Thus, represents a statement that is assumed to be unreachable.333Similar to __builtin_unreachable() from gcc. Consider all statements () in that are not covered by any path , then we can construct by replacing all such () with . We can represent this idea using the following rewrite rule:444We use the standard notation () to mean that any term matching the is rewritten to the term matching .
We may also apply the following rules to further simplify the resulting sub-program:
These rules preserve the (Truncation) property while also reducing the size (token count) of the resulting , which is beneficial for LLM prompting.
Slicing. We can further reduce the size of the truncated sub-program using program slicing (Weiser, 1981). We define a slice () to be a sub-program derived from by deleting (i.e., replacing by ) any statement () in that does not violate the condition:
Our main result is as follows. If () holds, then:
() holds by (Slice); and
() for all holds by (1) and (Generalization).
In other words, the single VC is sufficient to dispose of the entire partition (). We can apply standard slicing methods, such as Weiser’s back slicing algorithm (Weiser, 1981) (with as the slice criterion), to construct from . The back slicing algorithm traverses the Control Flow Graph (CFG), and deletes any statement that is not data- or control-flow-dependent on . The slicing algorithm is illustrated in Algorithm 2.
Example 3.0 (Slice-based Symbolic Execution).
We consider Example 3.1 once more. Here, the truncated slice () in Figure 2 (d) is a generalization of the infinite partition () representing all feasible paths through the -loop, including the Figure 2 (b) path (). An LLM can be used to verify the generalized verification condition (VC) encoded in natural language.
“Given the pre-condition and the code , does the post-condition hold?”
Path-based symbolic execution will infinitely unroll the -loop, generating a new VC for each loop iteration. In contrast, slice-based symbolic execution requires only a single VC to be checked. Furthermore, (, 5 lines, 28 tokens) is simpler and more compact than all of (, 10 lines, 56 tokens), (, 16 lines, 85 tokens), and (, 8 lines, 48 tokens). Truncation and slicing are useful for removing irrelevant statements while preserving paths, meaning that the corresponding prompt is simpler and more concise, thereby improving the accuracy of LLMs. ∎
Input: A CFG sub-graph ; reverse topological order Output: A 1 2 3Function GenSlice(, ): 4 ; getVars() 5 for do 6 if modifies any , or 7 conditional reads any then 8 9 10 11 12 end for 13 return 14 15 16 171ex 181ex Algorithm 2 Basic back-slicing algorithm.
Discussion. Our approach is related to path merging and loop invariants in traditional symbolic execution. Here, given a set of path constraints, represented as translated logical formulae , the idea is to find a formula that is generalization (). Thus, the verification conditions () can be combined into a single verification condition (), helping to mitigate the path explosion problem. Similarly, the (possibly infinite) set of path constraints through a loop can be generalized into a loop invariant , such that (), allowing the loop to be handled without infinite unrolling. However, loop-invariant discovery over formulae is difficult and undecidable in the general case. In contrast, our approach avoids the problem, since the -constraints are never translated into a different representation.
3.3.2. Partitioning
Input: The Control-Flow Graph (CFG) Output: A coverage-based partitioning Globals : Coverage map (), end node () 1 2 3Function GenPartitions(): 4 if then return 5 6 7 8 if then return 9 10 for do 11 12 GenPartitions() 13 return 14 151ex Algorithm 3 Path partitioning algorithm
Section 3.3.1 describes the construction of a truncated slice given a partition () that subset of paths (). In principle, any partitioning () can be used. However, a good partitioning aims to minimize the slice () size for each , , in order to simplify the prompts ultimately sent to the LLM. Our approach is to construct a partitioning based on path coverage.
Path Coverage. Under Section 3.1, we define a path () to be a linear program over statements or branches (represented as assertions) from . Here, we define a Control Flow Graph (CFG) path to be a sequence of nodes (a.k.a. locations) through the CFG representation of . Given a CFG path (), we define path coverage to be the set (), i.e., the reinterpretation of () from a sequence to a set. It is common for distinct paths to have the same coverage; for example, different iterations of the same loop have distinct sequences, but produce equivalent sets (i.e., covering the same nodes). We make two key insights. First, paths with distinct coverage will have distinct truncations, since each path must differ by at least one location. This will be used as the basis for partitioning. Secondly, the number of distinct paths w.r.t. coverage is finite, meaning that LLM-based symbolic execution (slice-based) necessarily terminates—even for unbounded loops.
Partitioning Algorithm. The core idea is to enumerate CFG paths () with distinct coverage (). Each partition () is implicitly defined as all paths with the same path coverage as (). The algorithm uses a Depth First Search (DFS) exploration over the CFG, as illustrated in Algorithm 3. Here, the algorithm takes a CFG representation of , and generates a partitioning as a set of representative paths () for each distinct coverage class (). The algorithm works as a standard DFS path construct algorithm, but also maintains a coverage map () that tracks previously-seen coverage prefixes. When a () is visited, the coverage map is consulted, and the path is pruned if the prefix has been observed before. The coverage map ensures that Algorithm 3 both (1) terminates, and (2) only returns paths (partitions) that differ by at least one location/node. Each output path () corresponds to a partition (), which is then used to construct a corresponding truncated sub-programs () and slice () that is used for LLM prompting.
Given two distinct coverages (, ), the resulting slices (i.e., , ) can sometimes be equivalent () if all distinguishing nodes are removed by the slicer. Furthermore, some truncated slices may generalize others, i.e., (). For example, under Algorithm 3, then is also a valid slice of (, Figure 2) representing zero iterations of the -loop. Nevertheless, our LLM-based symbolic execution algorithm orders VCs by size, in order to find a least refutation of the post-condition ().
Discussion. Algorithm 3 is guaranteed to terminate, since the number of possible distinct coverage sets is finite. Thus, unlike path-based symbolic execution, slice-based symbolic execution will always terminate, even for programs with unbounded loops. For example, the truncated slice () from Figure 2 (b) generalizes infinitely many paths through the loop (for any number of iterations). Furthermore, the size of each slice is always bounded by the size of the original program , whereas unrolled paths can exceed any length.
4. LLM-Based Symbolic Execution in Practice
The truncated slice-based symbolic execution algorithm of Section 3 is defined for an idealized programming language (the language of Hoare triples).
In this section, our aim is to port our abstract approach to real-world programming languages (e.g., C
/Java/Python).
Our first main insight is that the main symbolic execution and slicing algorithms are agnostic to the target programming language—provided there is a method for parsing the source code into some suitable representation for path generation and slicing.
Our second main insight is that, since LLMs are approximate oracles, the greater workflow can also be approximate—meaning that each individual step need not be perfectly precise,
provided the overall accuracy is not significantly degraded.
These insights significantly simplify our overall design.
An example of the workflow is illustrated in Figure 3.
First, our workflow uses a lightweight parsing framework (specifically, tree-sitter
(Brunsfeld and et al., 2025)) to parse the source code (a)
into an Abstract Syntax Tree (AST) representation (b).
From the AST, a Control-Flow Graph (CFG) is constructed, allowing for the finite enumeration of all path partitions using Algorithm 3.
For each enumerated partition, a truncation and slice is constructed (c), by removing statements that are not control- or data-flow dependent on the post-condition , using Algorithm 2.
Next, the slice is rendered back into the original source language (d).
Essentially, this means emitting sliced path elements, deleting (i.e., not emitting) non-sliced path elements, and replacing (i.e., replace by assume(0)
) all other non-path elements.
The resulting slice generalizes a (possibly infinite) set of paths through the program.
Finally, the slice is used to construct a text-based prompt (e) used to query an LLM (f).
steps (c)-(f) are repeated for each partition using Algorithm 1, until exhaustion or the discovery of a counterexample.
4.1. Partitioning and Truncated Slice Generation
Parsing.
We use the tree-sitter
framework (Brunsfeld and et al., 2025) to parse the source code into an AST.
Each AST node comprises a node type representing some syntactic element, the file range (source file and offset range) from which the element was parsed, and zero or more children of sub-elements.
The AST is unified so that the shared language features are mapped to the same representation, as illustrated in Figure 4.
Although tree-sitter
does not guarantee perfect parsing, this is tolerable under our approximate design.
Our approach also avoids heavyweight and specialized compiler frameworks necessary for precise parsing (e.g., clang
(Lattner and Adve, 2004) for C
, javac
(Arnold et al., 2005) for Java, and CPython (van Rossum, 1995) for Python), significantly simplifying the implementation.
CFG Construction. The next step is to lower the AST into a Control Flow Graph (CFG), where each node represents a statement or a condition, followed by one (or more) control-flow edges to successor nodes. The CFG is language agnostic, and common language features (e.g., if-then-else, while-loops, etc.) are lowered into common CFG patterns.
Slice generation.
The path partitioning is generated using Algorithm 3.
For each partition , a corresponding truncated sub-program is generated by substituting non-covered nodes/sub-programs with an assume(0)
statement, followed by simplification.
From this, a truncated slice is generated by applying Weiser’s back-slicing algorithm (as illustrated in Algorithm 2).
Here, the variables from the post-condition () are used as the slicing criterion, and Algorithm 2 removes all nodes from of which is not control- or data-flow dependent.
4.2. Truncated Slice Rendering
So far, each slice () is represented as a sub-graph of the (truncated) CFG. Our ultimate target is an off-the-shelf LLM, meaning that the slice must be rendered back into a generic text form. For this, we leverage the underlying AST representation based on file ranges. For the rendering algorithm, we use an interval tree () mapping ranges to strings. For each CFG node (and the corresponding AST statement ), we insert all into .
Context.
We also include the AST path from the root down to each sliced (), such as the enclosing function, control-flow context (if-then-else, while-loops, etc.), or class definitions (for Java).
This ensures that the rendered slice is coherent (reparsable) code, rather than a collection of isolated statements.
Furthermore, it is often useful to include all relevant non-executable AST nodes from which each depends, including
variable declarations, type declarations, member declarations, function declarations, global declarations, macro definitions, etc.
Such dependencies are matched based on the AST name.
For example, given the executable statement (xs = insert(xs, x)
), then any function declaration or macro definition matching the name “insert
” can be included in the context.
Since we rely on lightweight parsing and not a compiler front-end with semantic information, name-based matching may over-approximate dependencies.
However, this is allowable under our approximate workflow, and the LLM will often ignore irrelevant context.
The slice renderer also preserves the original formatting, including the preceding and succeeding whitespace and comments.
Preserving such information is not strictly necessary for C
and Java, but is essential for whitespace-sensitive languages such as Python.
Furthermore, source-code comments can provide additional context, such as the intent of programmers, which can assist LLM inference.
Example.
An example rendered slice is illustrated in Figure 3 (d).
Here, all sliced executable statements are preserved, as well as the enclosing context (e.g., function declaration for f()
Furthermore, the inner if-then-else has been simplified to an assertion, and the formatting (whitespace) has been preserved.
The resulting slice is coherent C
code in its own right.
Crucially, each path through the slice corresponds to an equivalent path in the original programming, meaning that any inference on the slice is also a valid inference for the original code.
LLM Inference.
Once the rendered slice has been generated, the final step is prompt construction and LLM inference.
This step can be highly customized,
but for our basic design,
we use a prompt structure that mirrors a Hoare triple where is a pre-condition, is a post-condition, and is the generated slice (see Example 2.3, Example 3.2, and Figure 3 (e)).
In addition, some basic instructions for the LLM are provided,
such as the output format.
Here, we assume that the LLM will generate one of two possible responses to the prompt, namely PASS
(the post-condition holds) or FAIL
(the post-condition does not hold).
4.3. Implementation
We have implemented our workflow in the form of the AutoBug tool.
AutoBug takes as input a program in a supported programming language (currently C
/Java/Python), pre- and post-conditions ( and ) expressed as code, constraints, or natural language.
AutoBug automatically decomposes the program into a sequence of truncated slices, and then constructs a sequence of prompts (a.k.a. verification conditions) to be sent to the LLM for inference.
The slices are ordered by size to find the least counterexample to the post-condition where applicable.
AutoBug is also language-agnostic, except for the parser and some elements of the renderer.
Furthermore, AutoBug is lightweight and approximate by design—significantly simplifying the implementation (i.e., it does not rely on language-specific compiler infrastructure).
This also reflects the nature of LLM inference, which is heuristic by nature rather than relying on precise parsing and semantic analysis, but can still make useful inferences for many real-world applications.
The implementation is also designed to be LLM-agnostic: since most modern LLMs expose a common API interface, AutoBug can be configured to query different models without changes to the core workflow.
5. Evaluation
To evaluate the effectiveness of LLM-based symbolic execution implemented by AutoBug, we answer the following research questions:
- RQ.1:
(Accuracy) What is the accuracy of AutoBug compared to baselines?
- RQ.2:
(Scalability) Can AutoBug scale to real-world large-scale programs?
- RQ.3:
(Language Agnosticism) Can AutoBug support multiple programming languages?
5.1. Experimental Setup
5.1.1. Dataset
We evaluated AutoBug on two widely-used datasets: REval and CodeForces. REval (Chen et al., 2024b) is designed to evaluate the code-reasoning abilities of LLMs. It consists of 85 Python solutions to LeetCode problems, some of which include accompanying test suites. The subjects in REval usually employ multiple control flow constructs, such as nested conditions and loops. This dataset was released in June 2024. In addition, we included the CodeForces dataset, which presents more challenging programming tasks. CodeForces is a programming contest platform that regularly publishes new problem sets. From this platform, we collected 662 subjects written in both Python and C
, published in June 2025. To the best of our knowledge, this is the most recent and largest open dataset currently available from the platform. The subjects in both datasets typically span around 100 lines of code, with token counts reaching up to 900.
To conduct the evaluation, we also needed to construct non-trivial Hoare triples for each subject. This presented a challenge, as both REval and CodeForces provide only problem descriptions in natural language, source-code solutions, and test suites in some cases. To generate Hoare triples, we employed three different strategies:
REval-Desc: We treated the program descriptions in natural language as the post-conditions, and used any restrictions on test input when available as the pre-conditions. This strategy was applied to the REval dataset to construct REval-Desc with the corresponding pre- and post-conditions.
CodeForces-Automatic: In this strategy, we used an LLM to automatically generate pre- and post-conditions in natural language based on the program descriptions from CodeForces. This resulted in the dataset CodeForces-Automatic for evaluation.
Mixed-Manual: Here, we manually annotated formal pre- and post-conditions in executable code based on implicit properties of the original problems, e.g., by adding assertions to verify the correctness of a sorting algorithm. Given the substantial manual effort required, we randomly sampled 30 subjects in both REval and CodeForces, covering both Python and
C
solutions, and applied this strategy to them. These formed another dataset Mixed-Manual.
5.1.2. Baselines
For comparison, we selected two main types of baselines. For the first baseline, we used an “ad-hoc” LLM-based program analysis over the entire subject without decomposition. This baseline queries the LLM for a counterexample (if one exists) or whether the post-condition always holds for inputs that satisfy the pre-condition. This baseline aims to show the effectiveness of the systematic partitioning provided by our tool AutoBug.
The second baseline includes traditional symbolic execution tools.
For Python subjects, we selected CrossHair (Schanely, 2017) as a baseline. Although there are other traditional symbolic execution tools for Python, such as PyExZ3 (Ball and Daniel, 2015) and pySym (Bann, 2016), they are no longer maintained (unmaintained for six years or more at the time of writing). For the C
subjects, we compared our tool with KLEE (Cadar et al., 2008), a well-established and mature traditional symbolic execution tool. A practical issue of these traditional tools is the lack of expressivity in the pre- and post-conditions, which must be specified in code or other formal languages. Therefore, we only compare AutoBug with traditional tools on the Mixed-Manual dataset, which contains conditions in the formal form of executable code.
5.1.3. Test Models
We selected five of the most popular openly available language models (at the time of writing)555https://ollama.com/search. Where possible, we also included different numbers of parameters to demonstrate the potential impact of the parameter number.
Meta’s Llama3 series (Grattafiori and et al., 2024): Instruction-tuned model optimized for general use cases such as dialogue and chatting, which was released in December 2024.
Microsoft’s Phi-4 (Abdin and et al., 2024): Model trained on high-quality data from filtered public domain websites and acquired academic books, designed to aid research on language models, which was released in December 2024.
DeepSeek’s DeepSeek-R1 (DeepSeek-AI, 2025): The first-generation reasoning model from DeepSeek, achieving performance comparable to OpenAI-o1 across math, code, and reasoning tasks with much smaller parameters, which was released in January 2025.
Google’s Gemma3 (Team, 2024): A lightweight family of multimodal models that features a larger context window of 128K, and released in March 2025.
Alibaba’s Qwen3 (Yang and et al., 2025): The latest version of the Qwen model family, designed to advance performance, efficiency, and multilingual capabilities, and it was released on April 2025.
All experiments were carried out using ollama
version v0.9.6. Each experiment was repeated three times, and we report the average results rounded to the nearest integer.
5.2. RQ.1: Accuracy
Method
To evaluate the accuracy of AutoBug in generating counterexamples and verifying post-conditions, we compared AutoBug with the “ad-hoc” LLM-based program analysis on REval-Desc, CodeForces-Automatic, and Mixed-Manual. In addition, we compared AutoBug with traditional symbolic execution tools CrossHair and KLEE on Mixed-Manual with formal pre- and post-conditions expressed in executable Python or C
code. We did not conduct the comparison on other datasets, since traditional tools cannot handle post-conditions in natural language. In each subject, we executed AutoBug and baselines, and reported their results. The accuracy in each subject was measured by comparing the result with the ground truth derived from the provided test suites, if available, or manual verification.
Main result
Table 2 shows the accuracy of AutoBug compared to the “ad-hoc” LLM-based program analysis across different datasets under different models. The experiments were conducted under the list of models (Model). For each dataset, we report the total number of subjects analyzed (Total), how many subjects were correctly analyzed by the tool (Correct), and the corresponding accuracy rate (Accuracy). Overall, AutoBug consistently achieves a higher accuracy than the baseline in all datasets using different forms of Hoare triples. On average, AutoBug improves accuracy from 84.7% to 90.6% on REval-Desc, from 79.2% to 86.4% on CodeForces-Automatic, and from 66.7% to 73.3% on Mixed-Manual. Notably, on these three (3) datasets under eight (8) language models, AutoBug outperforms the baseline in 20 out of 24 experiments (i.e., 83.3% scenarios), with the only exceptions on REval-Desc and Mixed-Manual when using the model Gemma3-27B. These results demonstrate the effectiveness of our approach in improving the accuracy of program analysis.
Model | Method | REval-Desc | CodeForces-Automatic | Mixed-Manual | ||||||
Total | Correct | Accuracy | Total | Correct | Accuracy | Total | Correct | Accuracy | ||
Llama3-8B | AutoBug | 85 | 82 | 96.5% | 662 | 654 | 98.8% | 30 | 23 | 76.7% |
Baseline | 85 | 79 | 92.9% | 662 | 653 | 98.6% | 30 | 22 | 73.3% | |
Llama3.1-8B | AutoBug | 85 | 84 | 98.8% | 662 | 616 | 93.1% | 30 | 23 | 76.7% |
Baseline | 85 | 71 | 83.5% | 662 | 547 | 82.6% | 30 | 18 | 60.0% | |
Llama3.3-70B | AutoBug | 85 | 72 | 84.7% | 662 | 617 | 93.2% | 30 | 26 | 86.7% |
Baseline | 85 | 70 | 82.4% | 662 | 589 | 89.0% | 30 | 21 | 70.0% | |
Phi-4-14B | AutoBug | 85 | 74 | 87.1% | 662 | 626 | 94.6% | 30 | 24 | 80.0% |
Baseline | 85 | 72 | 84.7% | 662 | 584 | 88.2% | 30 | 24 | 80.0% | |
DeepSeek-R1-70B | AutoBug | 85 | 85 | 100.0% | 662 | 561 | 84.7% | 30 | 19 | 63.3% |
Baseline | 85 | 82 | 96.5% | 662 | 508 | 76.7% | 30 | 17 | 56.7% | |
Gemma3-4B | AutoBug | 85 | 73 | 85.9% | 662 | 305 | 46.1% | 30 | 20 | 66.7% |
Baseline | 85 | 64 | 75.3% | 662 | 265 | 40.0% | 30 | 16 | 53.3% | |
Gemma3-27B | AutoBug | 85 | 72 | 84.7% | 662 | 598 | 90.3% | 30 | 20 | 66.7% |
Baseline | 85 | 76 | 89.4% | 662 | 547 | 82.6% | 30 | 23 | 76.7% | |
Qwen3-32B | AutoBug | 85 | 71 | 83.5% | 662 | 601 | 90.8% | 30 | 21 | 70.0% |
Baseline | 85 | 61 | 71.8% | 662 | 499 | 75.4% | 30 | 21 | 70.0% | |
Average | AutoBug | 85 | 77 | 90.6% | 662 | 572 | 86.4% | 30 | 22 | 73.3% |
Baseline | 85 | 72 | 84.7% | 662 | 524 | 79.2% | 30 | 20 | 66.7% |
Method | Mixed-Manual | ||
Total | Correct | Accuracy | |
AutoBug | 30 | 22 | 73.3% |
Baseline | 30 | 14 | 46.7% |
The comparison results with traditional symbolic execution tools are presented in Table 3, which shows the accuracy of AutoBug, and CrossHair and KLEE on Mixed-Manual with formal post-conditions in the form of executable code. We can see that AutoBug significantly outperforms these traditional tools. AutoBug successfully analyzes 73.3% of subjects, compared to just 46.7% for traditional tools. This performance gap is due to the inherent limitations of traditional symbolic execution (as illustrated in Section 2.2), such as the difficulties in reasoning about nested loops and complex constructs. Moreover, compared to traditional tools, AutoBug also demonstrates greater applicability by supporting expressive post-conditions, including those written in natural language.
Impact of different LLMs
We evaluated AutoBug on multiple different LLMs, even with different numbers of parameters as shown in Table 2: Meta’s Llama3, Llama3.1 and Llama3.3 models with 8B and 70B parameters, Microsoft’s 14B Phi-4 model, the 70B distilled version of DeepSeek-R1 model, Google’s Gemma3 model with 4B and 27B parameters, and Alibaba’s 32B Qwen3.
It is interesting to note that the accuracy varies significantly under different models, even with different model sizes. Among the models used, the most significant accuracy improvement stems from Llama3.1-8B, where AutoBug achieves 98.8% accuracy compared to the baseline’s 83.5%. This highlights the benefits of partitioning in improving the accuracy of LLM reasoning for comparatively small models. In contrast, on larger models such as Llama3.3-70B and DeepSeek-R1-70B, such improvements in accuracy are more subtle, mainly due to the fact that these larger models inherently have a stronger reasoning ability. However, the result is consistently positive for larger models, with the exception of Gemma3-27B. This is a positive result—smaller LLMs can be run locally on consumer hardware, and AutoBug exhibits a clear benefit for these use cases.
Under Gemma3-27B, AutoBug outperforms the baseline on CodeForces-Automatic, but shows weaknesses on REval-Desc and Mixed-Manual. Upon closer investigation, we attribute this to Gemma3-27B’s tendency to overinterpret problem specifications and penalize even minor mismatches. For example, during slicing, AutoBug may omit certain input-handling statements if they are irrelevant to the target context, resulting in incomplete input-processing logic. Gemma3-32B penalizes such incompleteness and reports unsatisfactory when inputs are not fully processed, while Gemma3-4B tends to overlook these issues and focuses on reasoning about the main task.
Prompt size
Table 4 shows the average token counts used in the LLM prompts for both the baseline and AutoBug across different datasets. To calculate the token count in each LLM prompt, we use the tool tiktoken
(Jain, 2025) with the corresponding Byte-Pair Encoding (BPE) token encoding, which is also used by GPT-4o mini (OpenAI, 2024). To clearly demonstrate the reduction ability of AutoBug, we also reported the reduction ratio achieved by AutoBug, compared with the baseline.
As shown in Table 4, AutoBug effectively reduces token counts—reducing the length and complexity of the prompt—with average reductions of approximately 9.5%, 25.3%, and 26.3% compared to the baseline for REval-Desc, CodeForces-Automatic, and Mixed-Manual, respectively.
Method | REval-Desc | CodeForces-Automatic | Mixed-Manual |
Baseline | |||
AutoBug | |||
Reduction Ratio | 9.5% | 25.3% | 26.3% |
Case study
To illustrate the advantages of AutoBug over the baselines, we use a concrete example in Figure 5. This example consists of a buggy Python function shown in Figure 5 (a), which is extracted from the REval dataset (Chen et al., 2024b). Here, the pre- and post-conditions are formally annotated with the corresponding PRE
and POST
comments. The pre-condition specifies that value is not empty, which is indicated by the (assume len(value) > 0
) statement located at the beginning of the function. The post-condition of the function states that . This condition does not always hold, since the given function rounds up any decimal input ending in “.5
”. For example, , which violates the post-condition.
We find that most LLMs cannot correctly reason over the whole program. A typical response from Llama3.1-8B (Grattafiori and et al., 2024) when fed with the entire original program is as follows:
“For inputs with a single decimal point, removing trailing zeros ensures that the absolute value of res (the rounded integer) is less than or equal to the absolute value of the original float ().”
The LLM then proceeds to incorrectly declare that the post-condition of the program will always be satisfied. In this case, the LLM is distracted over the trailing zero removal branch so much, thereby it does not actually try to analyze what the program is supposed to do with the converted value—i.e., an example of hallucination.
(a) original Python function
(b) truncated slice #1
(c) truncated slice #2
In contrast, AutoBug decomposes the input program into truncated slices. Two of the generated slices from the original Python function are shown in Figure 5 (b) and (c). As illustrated in this example, we can see that AutoBug significantly reduces the complexity and size of the resulting truncated slices. In Figure 5 (b), only the elif
branch is considered, with the others truncated, forming implicit pre-conditions in the form of assume
-statements.
In Figure 5 (c), only the second if
branch is assumed to be taken. In addition, the original Python code (a) contains 430 tokens, which is reduced to 103 tokens ( reduction) for (b) and 69 tokens ( reduction) for (c).
Together with the reduced CFG complexity, the slices form a smaller and more targeted prompt, increasing the chance of a correct validation result.
Using these truncated slices, the LLM correctly infers that the post-condition holds for slice (b), but does not hold for slice (c).
For example, Llama3.1-8B outputs the following for (c) (emphasis original):
“In other words, given an input that satisfies the precondition (but has a decimal point followed by
’.5’
), the postcondition will not be satisfied. The result ofceil
will exceed the absolute value of the original float. Therefore, we conclude that the postcondition indicated by theassert
statement with aPOST
comment is not always satisfiable.”
Since there exists a counterexample for (c), AutoBug has shown that the post-condition does not hold for the original program (a).
We also attempted to analyze Figure 5 (a) using CrossHair, a traditional symbolic execution engine for Python. However, CrossHair terminates after failing to find any violation of the post-condition. We believe that this is due to an incomplete search or incomplete solving of the path constraint by the underlying SMT solver (z3). In contrast, AutoBug is capable of exhaustively enumerating all path partitions, and the underlying LLM correctly infers the corresponding slice. This case study, although it involves a relatively simple program, highlights the limitations of traditional methods when compared to AutoBug.
5.3. RQ.2: Scalability
Method
To evaluate the scalability of AutoBug, we selected eight (8) real-world large-scale programs: X11, ncurses, nano, cURL, SQLite3, OpenSSH, FFmpeg, and Zstd, which are widely-used benchmarks for evaluating the performance of bug-finding techniques. Table 5 summarizes the statistics of these subject programs, including the lines of code (#LoC) and the token counts (#Token). These subjects have lines of code up to 250K and have 197K–4960K token counts. In this experiment, we examined whether AutoBug could successfully analyze the bugs recently found by existing bug detectors (Meng et al., 2024b). These bugs were found in several X11 client applications (xinput, xlsclients, xmodmap, xset, xwininfo), as well as in ncurses, SQLite3, and other common UNIX libraries, respectively. To construct post-conditions, we negated the crash condition, such as () for NULL-pointer dereference. To demonstrate the effectiveness of AutoBug’s decompositions, we evaluated three forms: only relevant source files included (File), only relevant functions included (Func), and relevant code sliced by AutoBug (Slice) along a path that exercises the bug.
Subject | X11 | ncurses | nano | cURL | SQLite3 | OpenSSH | FFmpeg | Zstd |
#LoC | 244K | 54K | 24K | 172K | 225K | 144K | 255K | 35K |
#Token | 4960K | 438K | 197K | 1379K | 2341K | 1427K | 2761K | 370K |
Subject | Bug | Token Count | Llama3.1-8B | Gemma3-4B | ||||||
File | Func | Slice | File | Func | Slice | File | Func | Slice | ||
xinput | NULL-pointer | 4281 | 1197 | 178 | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ |
xlsclients | NULL-pointer | 140120 | 1090 | 229 | ✗ | ✗ | ✗ | ✗ | ||
xmodmap | NULL-pointer | 18683 | 2189 | 279 | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ |
xset | Divide-by-zero | 17078 | 3164 | 506 | ✗ | ✓ | ✓ | ✗ | ✓ | ✓ |
xwininfo | Buffer-overflow | 152020 | 2560 | 597 | ✗ | ✗ | ✗ | ✗ | ||
ncurses | NULL-pointer | 1446 | 271 | 201 | ✓ | ✓ | ✓ | ✗ | ✓ | ✓ |
nano | Environment race | 6716 | 698 | 125 | ✗ | ✓ | ✓ | ✗ | ✗ | ✓ |
cURL | assert-trigger | 21285 | 5581 | 202 | ✗ | ✓ | ✓ | ✗ | ✗ | ✓ |
SQLite3 | NULL-pointer | 2341085 | 213088 | 4691 | ✗ | ✗ | ||||
OpenSSH | assert-trigger | 10164 | 446 | 208 | ✗ | ✓ | ✓ | ✗ | ✓ | ✓ |
FFmpeg | Buffer-overflow | 118456 | 19242 | 4370 | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ |
Zstd | NULL-pointer | 60813 | 936 | 328 | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ |
Accuracy | — | — | — | — | 8.3% | 41.7% | 75.0% | 0% | 25.0% | 75.0% |
Results
The results are shown in Table 6. AutoBug demonstrates strong scalability to real-world, large-scale programs, with token counts for all cases remaining within the context window of LLMs, making LLM-based symbolic execution feasible. Moreover, AutoBug achieves high accuracy, successfully analyzing 75.0% of the bugs. In contrast, traditional symbolic execution tools suffer from the well-known path-exploration problem and fail to scale to these programs. Additionally, the token counts for these subjects are too large for the ad-hoc LLM-based analysis within a single prompt, making decomposition necessary.
Compared to naïve decomposition methods based on files and functions, AutoBug’s slice-based decomposition method is significantly more effective. Using slicing, AutoBug successfully analyzes 75.0% of the bugs, while the file-based and function-based methods achieve only 8.3% and 41.7% success rates, respectively. To further show the effectiveness of AutoBug’s slicing method, we examine the bug in xinput. The bug resides in a program that contains 244K sLOC (4.96M tokens), but the sliced code is reduced to just 178 tokens, as shown in Figure 6. With the code size and complexity dramatically reduced, the bug becomes apparent: the XIQueryDevice()
function returns NULL
(derived from an error-handling path), and this value is immediately passed to XIFreeDeviceInfo()
(the intermediate code was removed by the slicer), thus violating the post-condition. After truncation and slicing, the violation is readily apparent even to small LLMs. In contrast, the other decompositions based on files and functions result in 4,281 and 1,197 tokens, respectively. They include large amounts of irrelevant information that obscure the bug and cause LLMs to miss it.
5.4. RQ.3: Language Agnosticism
Method
AutoBug implements a lightweight workflow that is language-agnostic. To evaluate AutoBug’s ability to analyze programs written in different programming languages, we conducted this experiment using the same set of subjects implemented in Python, C
, and Java, respectively. Specifically, we use GPT-4o-mini to automatically translate the REval-Desc dataset (originally written in Python) into equivalent C
and Java versions. These translated programs form the C-REval-Desc and Java-REval-Desc datasets, each consisting of 85 programs. We evaluated the three datasets using the same LLMs and framework as in RQ.1, and then reported the average results.
Model | Method | Python-REval-Desc | C-REval-Desc | Java-REval-Desc | ||||||
Total | Correct | Accuracy | Total | Correct | Accuracy | Total | Correct | Accuracy | ||
Average | AutoBug | 85 | 77 | 90.6% | 85 | 74 | 87.1% | 85 | 76 | 89.4% |
Baseline | 85 | 72 | 84.7% | 85 | 67 | 78.8% | 85 | 71 | 83.5% |
Results
The results are illustrated in Table 7. Across these three datasets implemented in different programming languages, AutoBug maintains similar performance and consistently improves the accuracy of the analysis. Specifically, compared to the baseline—the ad-hoc LLM-based analysis, AutoBug increases the accuracy of the analysis from 84.7% to 90.6% in Python, 78.8% to 87.1% in C
, and 83.5% to 89.4% in Java, respectively. These results demonstrate the versatility of AutoBug in handling multi-language programs, which is the direct result of its language-agnostic workflow and the inherent ability of LLM to understand multiple languages. Notably, these improvements are achieved without reliance on precise parsers and/or language-specific compiler front-ends.
To evaluate traditional symbolic execution tools such as CrossHair and KLEE, annotating formal post-conditions requires significant manual effort. In RQ.1 (Section 5.2), we manually annotated formal post-conditions and have shown the effectiveness of AutoBug in the programs with formal post-conditions compared to traditional symbolic execution. In this experiment, we explored the potential of using LLMs to automatically generate post-conditions based on specifications and translate them into code suitable for symbolic execution. Specifically, for CrossHair, we instructed LLM to generate the PEP 315-suitable contract specification, while we prompted LLM to generate suitable calls to klee_make_symbolic
for KLEE. However, the results were far from satisfactory. The LLMs consistently failed to translate natural language post-condition specifications into concrete, executable code. The generated code frequently contains compiler or runtime errors, including incomplete implementations, missing function bodies, or absent header and import statements. For example, CrossHair was only able to analyze fewer than 10% of the translated dataset (8 out of 85 subjects). These findings indicate that substantial manual effort is still necessary when using traditional symbolic execution tools.
5.5. Discussion
Our results show that path-based decomposition of program analysis tasks is effective at improving LLM-based program analysis, especially for small LLMs and real-world problems. The significance is that it allows for higher accuracy to be achieved with smaller models that can be run on consumer-grade hardware. In addition, we prove that a lightweight and language agnostic workflow is feasible and can still achieve good results.
Limitations
LLM-based program analysis is applicable to program analysis tasks that cannot be handled by traditional means. That said, the approximate reasoning of LLMs is not suitable for all applications, even with improved accuracy. LLMs are also unlikely to ever achieve the same accuracy as traditional solvers for some tasks, such as solving systems of linear equations. As such, we propose LLM-based symbolic execution as a complementary method that does not necessarily replace traditional approaches for all use cases. Another limitation is that path-based decomposition may still explode, even when our approach is guaranteed to terminate. This is an inherent limitation of path-based reasoning. However, we believe that the decomposition based on the truncated slice is a significant mitigation.
Threat to validity
A primary concern is the potential for data contamination in LLMs (Sallou et al., 2024), where our evaluation datasets may have been included in the models’ training data. To mitigate this issue, we selected the subjects from CodeForces that were released in June 2025, after the models’ pre-training cutoff date. As shown in Table 2, AutoBug consistently outperforms the baseline on these subjects. This suggests that the performance of AutoBug is not a result of data contamination.
6. Related Work
Static program analysis via symbolic execution
Symbolic execution (King, 1976) is an established method for static programs whose origins also relate to early work formalizing programs as mathematical logic.
The idea is to execute symbolic states, representing sets of concrete inputs, allowing for the exhaustive exploration of program behavior.
Over the decades, many different symbolic execution engines and frameworks have been developed, such as:
KLEE (Cadar et al., 2008),
Owi (Andrès et al., 2024),
Symbolic PathFinder (SPF) (Păsăreanu and Rungta, 2010),
Java PathFinder (JPF) (Visser et al., 2004),
CrossHair (Schanely, 2017),
Angr (Shoshitaishvili et al., 2016),
S2E (Chipounov et al., 2012),
PyExZ3 (Ball and Daniel, 2015), etc.
Unlike our approach, such traditional tools translate paths into some underlying formal language for theorem proving, thus inheriting many of the limitations discussed in this paper.
Furthermore, most existing tools are specialized to a specific language (C
, Java, binary, etc.) and are closely integrated into specific compiler frameworks (e.g., LLVM (Lattner and Adve, 2004) for KLEE).
That said, LLMs use a fundamentally different type of reasoning compared to the deductive reasoning of theorem provers.
As such, traditional approaches are suitable for problems that can be handled by traditional methods and for applications where perfect accuracy is required.
LLM-based program analysis
One recent alternative to traditional program analysis methods is Large Language Models (LLMs). LLMs are very general tools and can be applied to a wide variety of tasks, including fuzzing (Asmita et al., 2024; Meng et al., 2024c), vulnerability detection (Zhou et al., 2024b), and program repair (Fan et al., 2023). Another recent innovation is LLM-based agents (Xi et al., 2025), which are algorithms where decisions are made by the LLM. Our core Algorithm 1 is traditional and not agent-based. However, an agentized version could be made, but decisions (e.g., which branch to explore first) could be deferred to the LLM.
Intersection between symbolic execution and LLMs
There are some other nascent works that combine LLMs and symbolic execution. HyLLfuzz (Meng et al., 2024a) focuses on improving concolic execution integrated in hybrid fuzzing. LLM-Sym (Wang et al., 2024) is an agent-based symbolic execution framework for Python code that uses an LLM to translate paths into traditional path constraints suitable for solving via z3 (de Moura and Bjørner, 2008). LLM-Sym is fundamentally different in that our approach avoids translation altogether, instead directly using the LLM itself as a solver. Since LLM-Sym still uses translation to z3, it inherits many of the limitations of traditional symbolic execution engines discussed in this paper. Similarly, Loopy (Kamath et al., 2023) aims to discover loop invariants using LLMs, which can then be applied to symbolic analysis. Our approach avoids the need for invariant discovery, since it is not based on translation.
7. Conclusion
In this paper, we introduce a variant of symbolic execution that uses an LLM as the underlying reasoning engine instead of a traditional theorem prover or SMT solver.
Our approach introduces a generic path constraint representation in terms of the original code—allowing the LLM to reason directly over the path constraint and avoiding translation into a (less expressive) formal language.
Our approach allows for a path-based decomposition of the analysis task into smaller (more tractable) subtasks, which use fewer tokens (helping scale), and are more targeted (helping accuracy).
We implemented our approach in the form of AutoBug—a practical LLM-based symbolic execution engine that supports multiple programming languages (i.e., language agnostic, supporting C
/Python/Java) without depending on heavyweight compiler infrastructure.
Our experimental results demonstrate measurable improvements in terms of both accuracy and scale, especially in smaller models that can run on consumer-grade GPUs.